### v1

# Towards Exploring Continual Fine-Tuning for Enhancing Language Ability in Large Language Models

Abstract
A common challenge towards the adaptability of Large Language Models (LLMs) is their ability to learn new languages over time without hampering the model’s performance on languages in which the model is already proficient (usually English). Continual fine-tuning (CFT) is the process of sequentially fine-tuning an LLM to enable the model to adapt to downstream tasks with varying data distributions and time shifts. This paper focuses on the language adaptability of LLMs through CFT. We study a two-phase CFT process in which an English-only end-to-end fine-tuned LLM from Phase 1 (predominantly Task Ability) is sequentially fine-tuned on a multilingual dataset – comprising task data in new languages – in Phase 2 (predominantly Language Ability). We observe that the “similarity” of Phase 2 tasks with Phase 1 determines the LLM’s adaptability. For similar phase-wise datasets, the LLM after Phase 2 does not show deterioration in task ability. In contrast, when the phase-wise datasets are not similar, the LLM’s task ability deteriorates. We test our hypothesis on the open-source Mistral-7B and LLaMA-3-8B models with multiple phase-wise dataset pairs. To overcome the deterioration, we propose two methods, based on layer freezing and generative replay. We show that our methods assist the LLM in improving its language ability while preserving its task ability.
1 Introduction
Figure 1: Comparing hidden activations for Mistral-7B and LLaMA-3-8B during our two-phase continual fine-tuning process. We prompt each model with examples from MTBench (zheng2024judging), and visualize the similarity between the mean hidden activations, for each model layer. For datasets that encode “similar” tasks (Alpaca & MultiAlpaca), model’s task ability does not decline (e.g., 3% gain for IFEval). For non-similar datasets (Instruct & MultiAlpaca), the task ability declines (e.g., 8% decline for IFEval). Here, Phase 2 model representations do not align with Phase 1’s; thus, suggesting greater model weight interference and a decline in task ability.
With ever-increasing adoption of LLMs in real world applications and expanding multilingual user bases of these applications, it is important to cater these models to wide enough multilingual audiences. Model training is compute hungry, and there is an abundance of both labelled and unlabelled data in English as compared to other languages (shaham2024multilingual). As such, it is imperative to find efficient ways to use pre-trained or fine-tuned models to improve performance on other languages. In this paper, we refer to a model’s ability in non-English languages as predominantly its language ability (LA), which can be achieved without relying on large amounts of data in those languages. Instead, we can exploit the predominantly task ability (TA) learned from English data.
To this end, researchers use techniques like continual pre-training, continual fine-tuning or language adaption to adapt models to a newer set of languages to enhance their language abilities (refer to §LABEL:sec::rw). While these techniques are effective, they are highly task-specific. Furthermore, existing techniques for multilingual LLMs rely on parallel data, old fine-tuning data, or old and new set of parameters. Parameter efficient techniques like LoRA (hu2022lora) are also widely used to efficiently fine-tune LLMs on multilingual data. However, such techniques show both: catastrophic forgetting on English and incapability to exploit the task ability that the model receives from the English fine-tuning data (aggarwal2024maple).
In such a setting, we want to enhance the model’s language ability (other than English) while preserving the task ability achieved via (firstly) English fine-tuning. This setting results in the challenge of catastrophic forgetting, i.e., the model’s task ability on English may decline while fine-tuning on multilingual data (mukhoti2023fine). Furthermore, a trivial solution that fine-tunes on the mixture of multilingual and English-only data may be sub-par (e.g., due to language relatedness (dhamecha2021role)). Hence, it is challenging to improve an LLM’s language ability while preserving its performance on English.
Our Approach.
We use a two-phase continual fine-tuning (CFT) technique for language adaption. We study the effects of various English and multilingual instruction tuning datasets when the models are fine-tuned in two phases: where Phase 1 is fine-tuning the model in English to improve its task ability and then fine-tuning it on a proportionally-sized multilingual dataset in Phase 2. In Phase 1, we use Alpaca (alpaca) and OpenOrca (OpenOrca), and in Phase 2 we use MultiAlpaca (multialpaca) and mOpenOrca (§3.1).
We perform this study on two open-source models, namely LLaMA-3-8B and Mistral-7B. We also use fine-tuned versions of them, namely LLaMA-3-8B-Instruct and Mistral-7B-Instruct, as off-the-shelf Phase 1 fine-tuned models. We quantify a model’s task ability based on its performance on four English datasets: (i) two for instruction following (i.e., IFEval (zhou2023instruction) and Alpaca Eval (li2023alpacaeval)) and (ii) two for reasoning tasks (i.e., MMLU (hendryckstest2021) and HellaSwag (zellers2019hellaswag)). Likewise, we quantify a model’s language ability based on its performance on (i) two question answering tasks (i.e., MLQA (lewis2019mlqa) and XQuAD (Artetxe:etal:2019)) and (ii) XLSUM (hasan-etal-2021-xl), a summarization task.
Our Contributions.
First, we observe that when phase-wise English and multilingual datasets encode different tasks, we see a decline in the Phase 2 model’s performance on English. On the other hand, when Phase 1 and Phase 2 datasets encode similar tasks, the Phase 2 model’s performance on English improves (refer to Figure 1). Second, to quantify the similarity of these phase-wise datasets, we introduce two metrics based on language-agnostic embeddings and model representations. We show that our quantification correlates with the decline in task ability (§3.3). Third, we study the efficacy of two tailored variants of existing CFT strategies to mitigate the decline in task ability after Phase 2 fine-tuning, while also boosting the language ability. The first strategy we study is generative replay, i.e., using instructions from a similar English counterpart of the Phase 2 dataset to generate replay data using the Phase 1 model. The second strategy uses heuristic-based layer freezing. Here, we use the weight difference between the Base and Phase 1 models to pick specific layers for freezing during Phase 2 fine-tuning. We study the gains in task and language ability of these strategies compared to specific baselines (§4).
2 Enhancing Language Ability through Continual Fine-tuning
A common recipe to training LLMs to learn new languages is to use a training paradigm that focuses on task and language adaption (chen2024improvinglanguageplasticitypretraining). Concretely, we define task adaption as the model’s ability to comprehend the input text and then provide a suitable output. We refer to language adaption as the model’s ability to perform those tasks in languages other than English.
In Task Adaption, the LLM is trained to follow instructions, usually using labeled English data. Language Adaption focuses on training the LLM to understand text from newer languages. Task adaptation leverages cross-lingual transfer, facilitating language adaptation to a certain degree. However, this process can result in a decline in the LLM’s task adaptation performance due to the risk of catastrophic forgetting. Despite this challenge, task adaptation often yields greater benefits compared to relying solely on cross-lingual transfer for language adaptation.
Continual Fine-tuning for Language Adaption.
To improve the language adaption of LLMs, we re-imagine the above recipe as a two-phase CFT process. We have:
  • Phase 1. We fine-tune a base LLM end-to-end on an English instruction dataset. Phase 1 aims to predominantly teach the LLM instruction following ability, which we refer to as task ability.
  • Phase 2. Here, we use the fine-tuned LLM from Phase 1 and further end-to-end fine-tune it on a Multilingual instruction dataset. Unlike chen2024improvinglanguageplasticitypretraining, in our setting, the data in Phase 2 is labeled. However, compared to Phase 1, Phase 2’s dataset is geared towards enhancing the LLM’s language ability, and comprises multiple languages with fewer data points per language.
This paper relates English fine-tuning with task ability enhancement as English fine-tuning predominantly helps in the task ability of LLMs. Whereas multilingual fine-tuning predominantly helps with an LLM’s language ability.
CFT for Language Adaption: Challenges.
The primary challenge in our two-phase fine-tuning process is that the LLM’s language ability must not come at the cost of its task ability. We impose two additional constraints based on real-world scenarios. First, in Phase 2, we cannot re-use Phase 1’s dataset. Often instruction fine-tuned LLMs are available without their corresponding datasets (e.g., Mistral-7B-Instruct (mistral)). Second, in Phase 2, we cannot use the weights of the Phase 1 model during training, as saving both old and new set of parameters on the GPU for training would be computationally expensive.
In a nutshell, we focus on CFT for language adaption for an LLM while preserving the model’s task ability.
Model Phase 1 (P1) Phase 2 (P2) IFEval ()($\uparrow$)( ↑ ) Alpaca Eval ()($\uparrow$)( ↑ ) MMLU ()($\uparrow$)( ↑ ) HellaSwag ()($\uparrow$)( ↑ ) Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca MultiAlpaca 0.364 0.395 0.12 0.16 0.552 0.573 0.581 0.616 0.404 0.436 Instruct 0.550 0.462 0.35 0.15 0.575 0.533 0.641 0.416 0.529 0.390 LLaMA-3-8B Alpaca 0.277 0.326 0.10 0.11 0.231 0.242 0.556 0.567 0.291 0.311 Instruct 0.735 0.182 0.14 0.10 0.340 0.239 0.533 0.278 0.437 0.2
Table 1: Task Ability results for two-phase Continual Fine-tuning (CFT). When the phase-wise datasets are similar (Definition 1 and Definition 2), task ability post Phase 2 (P2) fine-tuning consistently improves (denoted with green). When the phase-wise datasets are not similar, we see a significant decline in task ability post Phase 2 (P2) fine-tuning (denote with red).
3 Evaluating Task & Language Ability for Multilingual CFT
3.1 Experiment Setup & Evaluation Tasks
Fine-tuning Models.
We continually fine-tune open-source Mistral-7B (mistral) and LLaMA-3-8B (llama) LLMs for language adaption.
Fine-tuning Datasets. For our phase-wise datasets, we use the open-source Alpaca (alpaca), MultiAlpaca (multialpaca), and OpenOrca (OpenOrca) datasets. Alpaca is a self-instruct English-only dataset. MultiAlpaca is a multilingual dataset created by translating Alpaca’s seed tasks to 11 languages and using GPT-3.5-Turbo for response collection. The languages are in equal proportions and are “French”, “Arabic”, “German”, “Spanish”, “Indonesian”, “Japanese”, “Korean”, “Portuguese”, “Russian”, “Thai”, and “Vietnamese”. The appendix (§A.2) describes OpenOrca and mOpenOrca.
Fine-tuning Technique. We perform full fine-tuning with bf16 precision to study the effects of full fine-tuning with multilingual data in Phase 2 and its effect on task ability. We also wish to exploit the benefits gained via complete fine-tuning of these models, which may not be possible with parameter efficient fine-tuning (aggarwal2024maple; panda2024lotteryticketadaptationmitigating). However, in §4, we propose a heuristic-based layer freezing strategy to mitigate forgetting of task ability in which we freeze some layers and fine-tune the rest. For our experiments, we use Axolotl222github.com/axolotl-ai-cloud/axolotl/, an open-source framework to fine-tune LLMs. We conducted our experiments on NVIDIA A100 GPUs with 80 GB RAM.
Evaluation Tasks. To quantify an LLM’s task ability, we evaluate Phase 1 and Phase 2 models on two instruction-following tasks (i) IFEval (zhou2023instruction) and (ii) Alpaca Eval (li2023alpacaeval), (iii) MMLU (hendryckstest2021) for problem-solving and (iv) HellaSwag (zellers2019hellaswag) for commonsense reasoning ability. To quantify an LLM’s language ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering (MLQA (lewis2019mlqa) & XQuAD (Artetxe:etal:2019)) and summarization (XLSUM (hasan-etal-2021-xl)). Further details on these tasks are available in the Appendix (§A.3).
To evaluate our models on TA and LA, we use LM-Evaluation-Harness333github.com/EleutherAI/lm-evaluation-harness, which is a unified framework for zero/few-shot evaluations of LLMs. For both task and language ability, we use zero-shot evaluation. For additional details on the training setup, code, and evaluation tasks, we refer the reader to the Appendix (§A).
3.2 Task and Language Ability Results
Model Phase 1 Phase 2 MLQA ()($\uparrow$)( ↑ ) XLSUM ()($\uparrow$)( ↑ ) XQuAD ()($\uparrow$)( ↑ ) Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B Alpaca MultiAlpaca 0.229 0.288 0.012 0.060 0.290 0.602 0.177 0.317 Instruct 0.246 0.307 0.012 0.033 0.351 0.436 0.203 0.259 LLaMA-3-8B Alpaca 0.438 0.597 0.033 0.034 0.586 0.737 0.352 0.456 Instruct 0.609 0.321 0.048 0.027 0.712 0.417 0.456 0.255
Table 2: Language Ability results for two-phase Continual Fine-tuning (CFT). With green, we denote an improvement in language ability post Phase 2 fine-tuning. Likewise, we denote a decline in language ability with red. For MLQA and XQUAD we use F1 abstractive score, while for XLSUM we use ROUGE Score.
We compare the task and language ability of Mistral-7B and LLaMA-3-8B continually fine-tuned models on different phase-wise datasets [^4] When it is clear from the context, we use “Instruct” to denote the dataset used in Phase 1 to instruction fine-tune Mistral-7B-Instruct or LLaMA-3-8B-Instruct.. Table 1 presents the results for task ability, while Table 2 presents the results for language ability. Table 2 reports the average score across languages. We provide language-specific scores in the Appendix (§B).
Results Discussion. From Table 1, we see that for phase-wise datasets like Instruct and MultiAlpaca, the performance of the Phase 2 models trained on them declines for English. This decline occurs when they are continually fine-tuned on multilingual data in Phase 2. However, we see a jump in Mistral-7B’s language ability from the results for the multilingual generative tasks (Table 2). These models fine-tuned on multilingual datasets show catastrophic forgetting in English. However, for phase-wise datasets like Alpaca followed by MultiAlpaca, we see that models trained on them do not show a decline in task ability (Table 1). We also see a gain in these models’ language ability (Table 2) [^5] LLaMA-3-8B Instruct MultiAlpaca shows deterioration in LA. We explain this behavior in §4.3..
Additional Ablations. In the Appendix (§B), we also present results for OpenOrca-mOpenOrca phase-wise datasets. For Mistral-7B, we observe that the average task ability of the Phase 2 model (over Phase 1’s Mistral-7B-OpenOrca) marginally declines: 0.487 from 0.504. Whereas, for Mistral-7B-Instruct, the average decline in task ability is significant: 0.376 from 0.529. Likewise, for LLaMA-3-8B, the average task ability for LLaMA-3-8B OpenOrca mOpenOrca sees an increase of 0.415 from 0.404. In contrast, with Instruct-mOpenOrca as the phase-wise datasets, the task ability significantly drops, from 0.437 to 0.173.
Observation. With Table 1, we see that our two-phase CFT setup for language adaption shows an interesting trend: for certain pairs of phase-wise datasets (e.g., Alpaca & MultiAlpaca), the LLM after Phase 2 sees an improvement in the task ability (computed on English evaluation tasks). We notice that phase-wise datasets like Alpaca and MultiAlpaca have the same seed prompts. Alternately, the two datasets encode the same tasks in different languages. We hypothesize an LLM fine-tuned on either of these datasets learns the same task ability, and therefore, the second phase of CFT leads to lesser interference in the representation space. That is, an LLM continually fine-tuned on Alpaca & MultiAlpaca preserves its task ability across phases. We next define two metrics that aim to quantify the task-specific similarity of two datasets.
3.3 Phase-wise Datasets: Similarity of Representations
Dataset Embedding Similarity (DES).
To quantify whether two datasets encode the same tasks, we define DES that computes a similarity score using the dot product of the average representations (embedding) generated by a language-agnostic model.
Definition 1 (Dataset Embedding Similarity (DES)).
Given a language-agnostic text embedding model $\Theta$, and any pair of datasets $D_{1}$ and $D_{2}$, let DES be the function $f_{\textsf{DES}}:D\times D\rightarrow[0,1]$
$f_{\textsf{DES}}(D_{1},D_{2};\Theta)=\langle\mathbf{E}_{\Theta}(D_{1}),\mathbf{E}_{\Theta}(D_{2})\rangle$ (1)
Here, $\mathbf{E}_{\Theta}(D_{i})\in\mathbb{R}^{d},~{}\forall i\in\{1,2\}$ is the normalized mean embedding across samples in $D_{i}$.
Higher the DES score, more similar the embedding, i.e., greater similarity between $D_{1}$ and $D_{2}$. For $\Theta$, we use the language-agnostic sentence-tokenizer LaBSE (feng2020languageagnostic). We compute DES by encoding 500 random samples from Alpaca, MultiAlpaca, OpenOrca, and mOpenOrca, and measure $f_{\textsf{DES}}$ for each pair.
Fixing Alpaca as the Phase 1 dataset $D_{1}$, when the Phase 2 dataset $D_{2}$ is MultiAlpaca, the DES score is 0.924 and 0.792 for mOpenOrca. When $D_{1}$ is OpenOrca, the DES score for mOpenOrca as $D_{2}$ is 0.953 and 0.774 when $D_{2}$ is MultiAlpaca. For dataset pairs with similar tasks, we see a high DES score and relatively low scores for datasets with different tasks. That is, DES captures the (pair-wise) variation in task abilities of these datasets.
Model Parameter Difference (MPD). Another method of quantifying the similarity of tasks for two datasets $D_{1}$ and $D_{2}$ is to compute the difference between the parameters of models $\Theta_{1}$ (fine-tuned on $D_{1}$) and $\Theta_{2}$ (fine-tuned on $D_{2}$). Geometrically, the difference of the parameters captures the representation shift of $\Theta_{2}$ in the space defined by $\Theta_{1}$. If $D_{1}$ $D_{2}$ encode the same tasks, the combined shift by $\Theta_{2}$ should be relatively lower, compared to the shift if $D_{1}$ $D_{2}$ encode different tasks. Formally,
Definition 2 (Model Parameter Difference (MPD)).
Given any two models $\Theta_{1}$ and $\Theta_{2}$ fine-tuned on self-instruct datasets $D_{1}$ and $D_{2}$ respectively, from the same base model $\Theta_{B}$, let MPD be the function $f_{\textsf{MPD}}:\Theta\times\Theta\rightarrow\mathbb{R}_{\geq 0}$ s.t.
$f_{\textsf{MPD}}(\Theta_{1},\Theta_{2};\Theta_{B})=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{w}(\Theta_{1,i})-\mathbf{w}(\Theta_{2,i})\|_{2}$ (2)
[…]
The smaller the MPD score, the closer the fine-tuned models are in the parameter space. Fixing Mistral-7B as the base model $\Theta_{B}$, and $D_{1}$ as MultiAlpaca, we vary $D_{2}$ as one of AlpacaOpenOrca, and mOpenOrca, and observe the corresponding MPD scores. We normalize the MPD scores with the maximum observed score across all three models for a fair comparison. With $D_{2}$ as Alpaca, the MPD score is $0.294$. For $D_{2}$ as Instruct, MPD is $1.0$ and 0.55 when $D_{2}$ is OpenOrca. These scores show a similar trend to DES: for Alpacaand MultiAlpaca, the scores are lower, highlighting the similarities in the datasets in the parameter space. We see relatively higher scores for the other pair of models, implying a difference in the dataset pairs.
3.4 Visualizing Decline in Task Ability
Setup.
To explain the effect of similar phase-wise data sets on the LLM task ability, we look at the model representations when parsing English (as the task ability is computed over English). We feed MTBench (zheng2024judging) to the models, an English prompt dataset for testing, and visualize the similarity between the mean hidden activations, for each model layer. For the analysis, given an LLM $\Theta$ with $l$ layers, let $X_{\Theta}\in\mathbb{R}^{l\times d}$ be the mean hidden activations, across $n$ samples from MTBench.
Figure 2: We see a greater change in the variation of the representations for non-similar datasets (e.g., Instruct & MultiAlpaca) compared to similar datasets (e.g., Alpaca & MultiAlpaca). Interestingly, for LLaMA-3-8B the change is large across layers and a magnitude higher than Mistral-7B. For Mistral-7B, we see the later layers showing the most change.
t-SNE Visualization. Figure 1 depicts t-SNEs (vanDerMaaten2008) for $X_{\textsc{Mistral-7B}}$ and $X_{\textsc{LLaMA-3-8B}}$ LLMs, continually fine-tuned on the phase-wise datasets Alpaca & MultiAlpaca and Instruct & MultiAlpaca. We observe that for similar phase-wise datasets, the model before and after Phase 2 produces similar hidden activations. Contrarily, for non-similar phase-wise datasets, the hidden activations form distinct clusters, implying separation between the phase-wise activations. That is, the model representations for non-similar phase-wise datasets are well-separated. The model representations during Phase 2 do not align with Phase 1 representations; thus, resulting in greater model weight interference leading to a decline in task ability.
Visualizing Variance in Model Representations. Figure 1 provides some intuition for the correlation between phase-wise datasets and decline in task ability. To further understand the layer-wise behavior of the hidden activations, similar to chang-etal-2022-geometry, we compute covariance matrices $\Sigma_{\Theta}$ for each $X_{\Theta}$. Intuitively, $\Sigma_{\Theta}$ captures the variance in different directions for representations of hidden activations for $\Theta$.
We first compute the mean centered activation matrix $\bar{X}_{\Theta}=X_{\Theta}-\mu_{\Theta}$, according to $\mu_{\Theta}\in\mathbb{R}^{d}$. Next, we derive $\Sigma_{\Theta}=\frac{1}{l-1}\cdot\bar{X}_{\Theta}^{T}\bar{X}_{\Theta}\in\mathbb{R}^{d\times d}$. To compare the layer-wise variance in representations, we compute the L2-Norm of the difference of the matrices $\Sigma_{\textsc{Mistral-7B}}$ (Figure 2 (left)) or $\Sigma_{\textsc{LLaMA-3-8B}}$ (Figure 2 (right)) when continually fine-tuned on Alpaca & MultiAlpaca (blue lines) or Instruct & MultiAlpaca (red lines).
From the figures, we see clear evidence of representational change, both in terms of the magnitude of the change and the subset of layers that show a greater change. For Mistral-7B, the Phase 2 model after CFT with Instruct & MultiAlpaca, shows 3 to 4 times more variation in its representations compared to the model with Alpaca & MultiAlpaca phase-wise datasets. This gap is significantly larger for LLaMA-3-8B.
CFT Setup Task Ability (TA) Language Ability (LA) Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Dataset Strategy ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) Mistral-7B 0.55 0.35 0.575 0.641 0.529 0.246 0.012 0.351 0.203 MultiAlpaca 0.462 0.15 0.533 0.416 0.390 0.307 0.033 0.436 0.259 LF_H1 0.456 0.03 0.497 0.598 0.395 0.176 0.016 0.215 0.136 LF_H2 0.364 0.12 0.364 0.504 0.338 0.213 0.014 0.442 0.223 GR_5 0.540 0.17 0.540 0.611 0.465 0.311 0.008 0.428 0.249 GR_10 0.567 0.12 0.567 0.594 0.462 0.213 0.007 0.427 0.215 LoRA 0.383 0.09 0.579 0.625 0.42 0.289 0.043 0.518 0.283 ER_10 0.593 0.08 0.580 0.635 0.599 0.249 0.008 0.398 0.218 LLaMA-3-8B 0.735 0.14 0.340 0.533 0.436 0.609 0.048 0.712 0.456 MultiAlpaca 0.182 0.10 0.239 0.278 0.217 0.321 0.030 0.417 0.256 LF_H1 0.303 0.0 0.231 0.275 0.202 0.368 0.037 0.505 0.303 LF_H2 0.380 0.06 0.485 0.525 0.373 0.400 0.038 0.505 0.314 GR_5 0.269 0.01 0.516 0.316 0.279 0.437 0.019 0.593 0.349 GR_10 0.264 0.12 0.229 0.250 0.228 0.254 0.009 0.314 0.192 LoRA 0.196 0.0 0.280 0.235 0.179 0.007 0.008 0.005 0.007 ER_10 0.420 0.02 0.603 0.561 0.420 0.434 0.025 0.53 0.330
Table 3: Task and Language ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10) and Layer Freezing (LF_H1 & LF_H2). We also use LoRA (hu2022lora) and ER_10 as two baseline strategies. Here, we perform Phase 2 fine-tuning with rank 64 and quantisation bfloat16 for LoRA. For ER_10, we use the English dataset used in GR_5 with original responses. The Phase 1 dataset is Instruct for each row. The first two rows for both Mistral-7B and LLaMA-3-8B provide numbers for Instruct and Instruct-MultiAlpaca (from Table 1 & Table 2).
4 Mitigating Strategies for Multilingual CFT
To mitigate the decline in task ability, we study two CFT techniques, Generative Replay (GR) and heuristic-based Layer Freezing (LF). In Generative Replay, we consider a new English data generation method motivated by the correlation between dataset similarity and task ability (§3.2). With heuristic-based Layer Freezing, we employ specific heuristics to find out the subset of layers to freeze in the model during Phase 2 fine-tuning.
4.1 Generative Replay
Typically, Generative Replay (GR) is a technique that generates data from past distributions to be used alongside new task data for the continual fine-tuning of a model on a new task (shin2017continuallearningdeepgenerative). However, from §3.2, we see that if the phase-wise datasets encode similar tasks, the decline in task ability is mitigated. Based on this observation, we use the Phase 1 model to generate responses, in English, from the English counterpart of the multilingual dataset used for training in Phase 2. This generated replay dataset acts as a bridge between the distributions of Phase 1 and Phase 2.
During Phase 2 fine-tuning, we include varying quantities of this generated data: specifically, 5% (GR_5) and 10% (GR_10), of the Phase 2 dataset. As a baseline, we also fine-tune the models with a similar sized subset of the English counterpart with original responses [^6] This dataset may not be available for all multilingual datasets eg. Aya (singh2024ayadatasetopenaccesscollection). In that case, instructions can always be translated to English but it is not always practical to translate responses. Hence, this baseline is the best-case scenario for our GR strategy.. We refer to this baseline as English Replay (ER_10).
4.2 Heuristic-based Layer Freezing
Model regularization is an effective technique to mitigate the drop in the previous task’s performance in continual learning (e.g., EWC (kirkpatrick2017overcoming)). However, this is computationally inefficient as it requires using both the old and new sets of parameters. Instead, we use Layer Freezing (LF), a relatively efficient technique for use as a ‘regularizer’ to preserve task ability during Phase 2. We consider the following two variations to select the set of layers to freeze:
  1. 1.
    LF_H1: freezing a random set of 10 layers of the model from Phase 1 to be fine-tuned in Phase 2.
  2. 2.
    LF_H2: freezing the top-10 layers that have changed the most during Phase 1 fine-tuning (e.g., Mistral-7B Base to Mistral-7B-Instruct). We select these layers separately for Key, Query, and Value, for each attention head.
We present our results in Table 3 for both GR and LF. Along with English Replay (ER), we define another baseline in which we use LoRA (hu2022lora) for continually fine-tuning in Phase 2.
4.3 Results Discussion
From Table 3, we see that GR and LF successfully mitigate the decline in task ability and also show gains in language ability. For instance, Mistral-7B with GR_5 achieves better performance in MLQA and XLSUM when fine-tuned with MultiAlpaca. We also close the gap with Mistral-7B-Instruct  on IFEval, Alpaca Eval, MMLU, and HellaSwag with our mitigation strategies.
(a) Mistral-7B
(b) LLaMA-3-8B
Figure 3: Demonstrating extent of cross-lingual transfer in Mistral-7B and LLaMA-3-8B on a parallel dataset prepared by subsampling Flores (costa2022no). We find that the English activation cluster for LLaMA-3-8B is separated from the multilingual cluster, compared to Mistral-7B.
LLaMA-3-8B Doesn’t Show Consistent Improvement with our Mitigations.
From Table 3, while both GR and LF improve on the baseline LLaMA-3-8B-Instruct MultiAlpaca, the gains in task and language ability are not comparable to LLaMA-3-8B-Instruct.
To understand this further, for GR, we investigate the cross-linguality difference between LLaMA-3-8B and Mistral-7B. Like Figure 1, we plot t-SNEs of the mean model activations for the Mistral-7B and LLaMA-3-8B base models on two parallel datasets, English and Multilingual. We create the parallel datasets by subsampling data from Flores (costa2022no). In Figure 3, we see that the English activation cluster for LLaMA-3-8B is separated out from multilingual cluster, compared to Mistral-7B. This suggests that GR may not be as effective when the model has less cross lingual ability. While for LF, we acknowledge that our method to identify the layers to freeze may not be the best and better methods to identify which layers to freeze can be a direction for future work.
Last, but not the least, we acknowledge that LLaMA-3-8B-Instruct seems to be a strong model even on multilingual benchmarks. Hence, it is also important to evaluate Phase 1 models on these benchmarks first and then decide if the Phase 2 fine-tuning step should be undertaken or not.
With regards to LLaMA-3-8B-Instruct MultiAlpaca LA results in Table 2, we believe that this is due to lack of cross-linguality in LLaMA-3-8B-Instruct and less data in MultiAlpaca which fails to cause sufficient representation drift to improve the model’s performance.
Forgetting with LoRA. For Mistral-7B-Instruct and LoRA fine-tuning, we see an increase in language ability but a decline in task ability. But the decline is not as much as full fine-tuning. For LLaMA-3-8B-Instruct and LoRA, there is a greater decline in both task and language ability. The decline is similar (or slightly lower) than the full fine-tuning scenario. These results show that LoRA also suffers from forgetting when used for continual fine-tuning.
Additional Results. In the Appendix (§C), we repeat the same experiment from §3.4 to quantify the representation change in the fine-tuned models using our mitigating strategies. We see a trend similar to Figure 2. That is, a decrease in the variation in the model activations, compared to the baseline model trained on Instruct and MultiAlpaca. The decrease is more pronounced for Mistral-7B compared to LLaMA-3-8B. In Appendix §C, we also present TA and LA results for the Instruct-mOpenOrca phase-wise datasets.
5 Conclusion & Future Work
In this paper, to the best of our knowledge, we present a first study on the influence of the similarity of phase-wise datasets on the task and language adaptability of LLMs through CFT. Through extensive experiments on the Mistral-7B and LLaMA-3-8B models, we show that when datasets are similar, task ability is preserved; otherwise, it declines. To mitigate this, we propose layer freezing and generative replay as mitigating strategies based on specific heuristics. Our results indicate that these strategies help improve task performance while not compromising on the LLM’s language adaptability.
Future Work.
Our results show that there is no one-size-fits-all strategy to mitigate decline in task ability, among the strategies discussed. Future work can explore developing other parameter-efficient regularization methods that address the current computational challenges with methods like EWC or forgetting due to LoRA. One can also explore analytical notions for task similarity in datasets.
Appendix A Training Details
A.1 Hyperparameters for Fine-tuning and Training Setup
Hyperparameter Value Learning Rate $1\times 10^{-6}$ Epochs 4 Global Batch size 16 Scheduler Cosine Warmup Linear Warmup Steps 10 Optimizer AdamW (loshchilov2019decoupled) Weight Decay 0
Table A1: Hyperparameters for continual fine-tuning
A.2 Fine-tuning Datasets
OpenOrca is an English-only self instruct dataset, created to best mimic the Orca dataset (mukherjee2023orca), which is not publicly available. To create the multilingual version of OpenOrca, namely mOpenOrca, we follow sphinx to generate selective translations for a subset of OpenOrca. The subset contains 50k samples from the OpenOrca dataset and we selectively translate them to 11 languages which are also in MultiAlpaca. In total, we generate 550k examples for all languages.
A.3 Evaluation Tasks
In this paper, we consider two sets of benchmarks to evaluate task and language ability. We explain them briefly next.
Task Ability (TA).
To quantify an LLM’s task ability, we evaluate Phase 1 and Phase 2 models on the following tasks:
  1. 1.
    IFEval (zhou2023instruction): Instruction-Following Evaluation (IFEval) asses the ability of an LLM to follow natural language instructions. It comprises 500 verifiable instructions (e.g., “mention the keyword AI 3 times”). We choose IFEval as the instructions are verifiable and also test an LLM’s context understanding.
  2. 2.
    Alpaca Eval (li2023alpacaeval): This is an LLM-based automatic evaluator for instruction following models, to measure task ability. Like aggarwal2024maple, we evaluate our CFT models against text-davinci-003 responses on 800 instructions and use GPT4 (gpt-4-32k) as the evaluator.
  3. 3.
    MMLU (hendryckstest2021): Massive Multitask Language Understanding (MMLU) is a benchmark to assess an LLM’s knowledge and problem-solving abilities. It includes 57 subjects across domains like STEM, or law, with 16k MCQs in total.
  4. 4.
    HellaSwag (zellers2019hellaswag): This is a popular benchmark to evaluate the commonsense reasoning ability of an LLM. HellaSwag’s test split contains 10k samples in total.
Language Ability (LA).
To quantify an LLM’s language ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering and summarization.
  • Question Answering: MLQA (lewis2019mlqa) contains 5k extractive question-answering instances in 7 languages. The XQuAD dataset (Artetxe:etal:2019) consists of a subset of 240 paragraphs and 1190 question-answer pairs across 11 languages.
  • Summarisation: XLSUM (hasan-etal-2021-xl) spans 45 languages, and we evaluate our models in Arabic, Chinese-Simplified, English, French, Hindi, Japanese, and Spanish.
Model Phase 1 (P1) Phase 2 (P2) IFEval ()($\uparrow$)( ↑ ) Alpaca Eval ()($\uparrow$)( ↑ ) MMLU ()($\uparrow$)( ↑ ) HellaSwag ()($\uparrow$)( ↑ ) Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.494 0.482 0.31 0.32 0.601 0.582 0.612 0.562 0.504 0.487 Instruct 0.550 0.426 0.35 0.06 0.575 0.507 0.641 0.509 0.529 0.376 LLaMA-3-8B OpenOrca 0.377 0.425 0.09 0.07 0.579 0.599 0.571 0.564 0.404 0.415 Instruct 0.735 0.205 0.14 0.0 0.340 0.236 0.533 0.250 0.437 0.173
Table A1: Task Ability results for two-phase Continual Fine-tuning (CFT). With green, we highlight an increase in a model’s task ability post P2 fine-tuning. Likewise, red highlights a decline in a model’s task ability.
Model Phase 1 Phase 2 MLQA ()($\uparrow$)( ↑ ) XLSUM ()($\uparrow$)( ↑ ) XQuAD ()($\uparrow$)( ↑ ) Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B OpenOrca mOpenOrca 0.435 0.36 0.007 0.008 0.556 0.643 0.332 0.337 Instruct 0.246 0.155 0.012 0.040 0.351 0.323 0.203 0.173 LLaMA-3-8B OpenOrca 0.401 0.453 0.017 0.006 0.499 0.531 0.306 0.330 Instruct 0.609 0.604 0.048 0.048 0.712 0.713 0.456 0.455
Table A2: Language Ability results for two-phase Continual Fine-tuning (CFT). With green, we highlight an increase in a model’s language ability post Phase 2 fine-tuning. Likewise, red highlights a decline in a model’s language ability.
Appendix B Evaluating Language Ability for Multilingual Continual Fine-tuning
Task Ability.
Table A1 present the task ability numbers of our ablations on the OpenOrca-mOpenOrcaand Instruct-mOpenOrcadatasets using Mistral-7B and LLaMA-3-8B models. When the datasets are pairwise not similar, i.e., Instruct-mOpenOrca, Mistral-7B shows a significant decline in the average task ability, from 0.529 in Phase 1 to 0.376 in Phase 2. Likewise, LLaMA-3-8B also experiences a decrease, dropping from 0.437 to 0.173 on average.
In contrast, when the pairwise datasets are similar, i.e., OpenOrca and mOpenOrca, Mistral-7B sees a marginal drop between the phases $(0.504\rightarrow 0.487)$, on average. LLaMA-3-8B’s performance sees an improvement in the average task ability, from 0.404 to 0.415.
Language Ability.
Table A2 tabulates the results for language ability. We see an improvement in the average language ability for the OpenOrca-mOpenOrca dataset pair, for both Mistral-7B and LLaMA-3-8B. For Instruct-mOpenOrca, with LLaMA-3-8B, the average language ability is virtually the same across tasks. However, for Mistral-7B, we see a slight drop in the average language ability, driven primarily due to a drop in performance for MLQA.
Furthermore, Table B3, Table B4, and Table B5 present the language-specific results for MLQA, XLSUM, and XQuAD, respectively.
Model Phase 1 Phase 2 MLQA Dataset Dataset Phase 1 Phase 2 ar de es hi vi zh ar de es hi vi zh Mistral-7B Alpaca MultiAlpaca 0.143 0.337 0.331 0.149 0.385 0.031 0.172 0.485 0.529 0.196 0.336 0.009 Instruct 0.113 0.440 0.395 0.088 0.369 0.073 0.228 0.456 0.529 0.279 0.327 0.0222 LLaMA-3-8B Alpaca 0.320 0.538 0.563 0.438 0.611 0.155 0.552 0.672 0.765 0.573 0.784 0.237 Instruct 0.549 0.701 0.769 0.624 0.788 0.192 0.316 0.453 0.526 0.137 0.464 0.028 Mistral-7B OpenOrca mOpenOrca 0.374 0.504 0.511 0.395 0.600 0.226 0.298 0.506 0.572 0.274 0.481 0.030 Instruct 0.113 0.440 0.395 0.088 0.369 0.073 0.115 0.253 0.213 0.088 0.222 0.038 LLaMA-3-8B OpenOrca 0.262 0.545 0.565 0.369 0.568 0.099 0.437 0.549 0.622 0.462 0.625 0.024 Instruct 0.320 0.538 0.563 0.438 0.611 0.155 0.554 0.701 0.771 0.625 0.787 0.188
Table B3: MLQA: Language Ability results for two-phase Continual Fine-tuning (CFT).
Model Phase 1 Phase 2 XLSUM Dataset Dataset Phase 1 Phase 2 Arabic Chinese_simplified french Hindi Japanese Spanish Arabic Chinese_simplified french Hindi Japanese Spanish Mistral-7B Alpaca MultiAlpaca 0.001 0.012 0.025 0.001 0.012 0.023 0.022 0.034 0.112 0.016 0.067 0.106 Instruct 0.001 0.005 0.028 0.001 0.009 0.025 0.016 0.015 0.060 0.010 0.040 0.056 LLaMA-3-8B Alpaca 0.005 0.015 0.071 0.003 0.037 0.067 0.003 0.018 0.073 0.002 0.041 0.070 Instruct 0.008 0.015 0.092 0.004 0.080 0.087 0.002 0.013 0.055 0.001 0.055 0.051 Mistral-7B OpenOrca mOpenOrca 0.001 0.010 0.014 0.001 0.007 0.009 0.001 0.006 0.018 0.001 0.008 0.016 Instruct 0.001 0.005 0.028 0.001 0.009 0.025 0.007 0.017 0.092 0.005 0.030 0.088 LLaMA-3-8B OpenOrca 0.000 0.003 0.061 0.000 0.004 0.035 0.000 0.003 0.016 0.001 0.000 0.013 Instruct 0.008 0.015 0.092 0.004 0.080 0.087 0.007 0.015 0.091 0.004 0.082 0.087
Table B4: XLSUM: Language Ability results for two-phase Continual Fine-tuning (CFT).
Model Phase 1 Phase 2 XQuAD Dataset Dataset Phase 1 Phase 2 ar de el es hi ro ru th tr vi zh ar de el es hi ro ru th tr vi zh Mistral-7B Alpaca MultiAlpaca 0.194 0.379 0.248 0.374 0.224 0.418 0.150 0.185 0.454 0.475 0.088 0.613 0.692 0.657 0.713 0.670 0.679 0.661 0.385 0.666 0.734 0.148 Instruct 0.166 0.568 0.260 0.510 0.173 0.508 0.336 0.210 0.460 0.502 0.168 0.369 0.612 0.253 0.634 0.450 0.553 0.555 0.180 0.532 0.566 0.089 LLaMA-3-8B Alpaca 0.393 0.689 0.529 0.735 0.644 0.723 0.538 0.398 0.671 0.748 0.376 0.676 0.850 0.710 0.893 0.740 0.817 0.726 0.526 0.770 0.884 0.519 Instruct 0.659 0.795 0.702 0.852 0.715 0.810 0.609 0.594 0.728 0.834 0.533 0.444 0.580 0.244 0.657 0.241 0.586 0.493 0.092 0.580 0.558 0.113 Mistral-7B OpenOrca mOpenOrca 0.001 0.010 0.014 0.001 0.007 0.009 0.001 0.006 0.018 0.001 0.008 0.639 0.832 0.570 0.847 0.601 0.776 0.771 0.366 0.734 0.820 0.113 Instruct 0.166 0.568 0.260 0.510 0.173 0.508 0.336 0.210 0.460 0.502 0.168 0.256 0.457 0.320 0.443 0.256 0.409 0.215 0.245 0.364 0.428 0.162 LLaMA-3-8B OpenOrca 0.505 0.642 0.587 0.711 0.604 0.634 0.651 0.290 0.699 0.685 0.104 0.639 0.832 0.570 0.847 0.601 0.776 0.771 0.366 0.734 0.820 0.113 Instruct 0.659 0.795 0.702 0.852 0.715 0.810 0.609 0.594 0.728 0.834 0.533 0.654 0.793 0.703 0.852 0.718 0.808 0.606 0.600 0.729 0.836 0.540
Table B5: XQuAD: Language Ability results for two-phase Continual Fine-tuning (CFT).
Appendix C Mitigating Strategies
Here, we first visualize the impact of our mitigating strategies on the variance in model representations. Next, we provide ablate our findings for the Instruct-mOpenOrca phase-wise datasets.
Figure C4: Visualizing Variance in Model Representations for Mistral-7B Mitigating Strategies: We see a decrease in the variance of model representations for models trained using our mitigation strategies compared to vanilla Phase 2 models (see Figure 2).
Visualizing Variance in Model Representations.
In Figure C4, we repeat the same experiment as in § 4.5 to quantify the representation change in the fine-tuned models using our mitigating strategies. The trend seen is expected from §4.5: we see a decrease in the variation in the model activations, compared to the baseline model trained on Instruct and MultiAlpaca.
For the mitigating strategies that are curated to curb representational change, i.e., LF_H2, GR_5, and GR_10, we see that the corresponding curves have lesser change than the baseline Phase 2 model, Mistral-7B Instruct MultiAlpaca. That is, there is less representational change for LF_H2, GR_5, and GR_10 compared to Mistral-7B Instruct MultiAlpaca.
Our generative replay techniques are the closest in the representational change to Mistral-7B Instruct. This ‘closeness’ also improves its task and language ability performance compared to the vanilla Phase 2 model, Mistral-7B Instruct MultiAlpaca(refer to Table 1 and Table 2).
Additional Ablations.
We also present the impact of our mitigating strategies for the Instruct-mOpenOrca phase-wise datasets on Mistral-7B. Table C6 presents these results.
We see that LF_H2 achieves moderate success, especially in maintaining the language ability for MLQA (0.258) and XQUAD (0.527). However, task ability shows some decline (e.g., IFEval (0.401) and Alpaca Eval (0.048)), compared to the baseline. Furthermore, GR_5 results in lower task ability (IFEval = 0.281), while GR_10 performs slightly better in task ability (e.g., MMLU = 0.483, HellaSwag = 0.494).
Among the baselines, ER_10 performs similarly to the generative replay strategies, with modest improvements in task ability (e.g., IFEval = 0.367, MMLU = 0.479), but still struggles in language ability. Perhaps LoRA shows the best overall performance among the strategies for maintaining task ability (e.g., IFEval = 0.587, MMLU = 0.567, HellaSwag = 0.591) with reasonable retention of language ability (e.g., XQUAD = 0.354).
Note.
These results show that no single strategy is perfect, and future work may need to combine these strategies or develop new approaches to address the balance between task and language ability retention across phases.
CFT Setup Task Ability Language Ability Model Phase 1 Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Dataset Dataset Strategy Mistral-7B Instruct 0.55 0.35 0.575 0.641 0.529 0.246 0.012 0.351 0.203 mOpenOrca 0.426 0.06 0.507 0.509 0.376 0.155 0.040 0.323 0.173 LF_H2 0.401 0.048 0.518 0.487 0.364 0.258 0.060 0.527 0.282 GR_5 0.281 0.027 0.478 0.495 0.320 0.167 0.042 0.305 0.171 GR_10 0.305 0.013 0.483 0.494 0.324 0.150 0.038 0.238 0.142 LoRA 0.587 0.13 0.567 0.591 0.469 0.167 0.027 0.354 0.183 ER_10 0.367 0.025 0.479 0.493 0.341 0.157 0.042 0.305 0.168
Table C6: Task and Language ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10) and Layer Freezing (LF_H1 & LF_H2). We also use LoRA (hu2022lora) and ER_10 as two baseline strategies. Here, we perform Phase 2 with rank 64 and bf16 for LoRA. For ER_10, we use the English dataset used in GR_5 with original responses.
Appendix D Resources Used
We used 4 NVIDIA A100 GPU (80 GB) with a 96 core AMD CPU to run our inferences. One Finetuning Run with MultiAlpaca took 4 hours while for mOpenOrca it took 12 hours.
the list of model and the URL with checkpoints available and licenses are listed below:
  • LLaMA-3-8B : meta-llama/Meta-Llama-3-8B License: llama3
  • Mistral-7B : https://huggingface.co/mistralai/Mistral-7B-v0.1 License: Apache-2.0

### v3

@@
-Towards Exploring Continual Fine-Tuning for Enhancing Language Ability in Large Language Models
+Exploring Two-Phase Continual Instruction Fine-tuning for Multilingual Adaptation in Large Language Models
 Abstract
-A common challenge towards the adaptability of Large Language Models (LLMs) is their ability to learn new languages over time without hampering the model’s performance on languages in which the model is already proficient (usually English).
-Continual fine-tuning (CFT) is the process of sequentially fine-tuning an LLM to enable the model to adapt to downstream tasks with varying data distributions and time shifts.
-This paper focuses on the language adaptability of LLMs through CFT.
-We study a two-phase CFT process in which an English-only end-to-end fine-tuned LLM from Phase 1 (predominantly Task Ability) is sequentially fine-tuned on a multilingual dataset – comprising task data in new languages – in Phase 2 (predominantly Language Ability).
-We observe that the “similarity” of Phase 2 tasks with Phase 1 determines the LLM’s adaptability.
-For similar phase-wise datasets, the LLM after Phase 2 does not show deterioration in task ability.
-In contrast, when the phase-wise datasets are not similar, the LLM’s task ability deteriorates.
-We test our hypothesis on the open-source Mistral-7B and LLaMA-3-8B models with multiple phase-wise dataset pairs.
-To overcome the deterioration, we propose two methods, based on layer freezing and generative replay.
-We show that our methods assist the LLM in improving its language ability while preserving its task ability.
+A key challenge for Large Language Models (LLMs) is improving their Multilingual instruction-following ability over time without deteriorating their ability in languages they already excel at, typically English.
+In this paper, we study a two-phase Continual Fine-tuning (CFT) setup toward improving a model’s Multilingual adaptability.
+Concretely, we consider a two-phase CFT process in which an English-only end-to-end instruction fine-tuned LLM (Phase 1) is sequentially fine-tuned on a multilingual instruction dataset (Phase 2).
+Across Mistral-7B and LLaMA-3-8B and multiple dataset pairs, we show that instructional similarity between phases is critical: aligned datasets preserve or improve English while boosting multilingual ability, whereas misaligned datasets cause English degradation.
+We show that this degradation arises from representation shift during CFT, and that targeted mitigation strategies, including generative replay and heuristic-based layer freezing, reduce this shift and improve multilingual adaptation.
+Exploring Two-Phase Continual Instruction Fine-tuning for Multilingual Adaptation in Large Language Models
+Divyanshu Aggarwalthanks: Joint First Author1, Sankarshan Damlefootnotemark: 1, Navin Goyal2, Satya Lokam1, Sunayana Sitaram2 1Microsoft  2Microsoft Research India {divyanshu.aggarwal,t-sandamle,navingo,satya.lokam,sunayana.sitaram}@microsoft.com
 1 Introduction
-Figure 1: Comparing hidden activations for Mistral-7B and LLaMA-3-8B during our two-phase continual fine-tuning process.
-We prompt each model with examples from MTBench (zheng2024judging), and visualize the similarity between the mean hidden activations, for each model layer.
-For datasets that encode “similar” tasks (Alpaca & MultiAlpaca), model’s task ability does not decline (e.g., 3% gain for IFEval).
-For non-similar datasets (Instruct & MultiAlpaca), the task ability declines (e.g., 8% decline for IFEval).
-Here, Phase 2 model representations do not align with Phase 1’s; thus, suggesting greater model weight interference and a decline in task ability.
-With ever-increasing adoption of LLMs in real world applications and expanding multilingual user bases of these applications, it is important to cater these models to wide enough multilingual audiences.
-Model training is compute hungry, and there is an abundance of both labelled and unlabelled data in English as compared to other languages (shaham2024multilingual).
-As such, it is imperative to find efficient ways to use pre-trained or fine-tuned models to improve performance on other languages.
-In this paper, we refer to a model’s ability in non-English languages as predominantly its language ability (LA), which can be achieved without relying on large amounts of data in those languages.
-Instead, we can exploit the predominantly task ability (TA) learned from English data.
-To this end, researchers use techniques like continual pre-training, continual fine-tuning or language adaption to adapt models to a newer set of languages to enhance their language abilities (refer to §LABEL:sec::rw).
-While these techniques are effective, they are highly task-specific.
-Furthermore, existing techniques for multilingual LLMs rely on parallel data, old fine-tuning data, or old and new set of parameters.
-Parameter efficient techniques like LoRA (hu2022lora) are also widely used to efficiently fine-tune LLMs on multilingual data.
-However, such techniques show both: catastrophic forgetting on English and incapability to exploit the task ability that the model receives from the English fine-tuning data (aggarwal2024maple).
-In such a setting, we want to enhance the model’s language ability (other than English) while preserving the task ability achieved via (firstly) English fine-tuning.
-This setting results in the challenge of catastrophic forgetting, i.e., the model’s task ability on English may decline while fine-tuning on multilingual data (mukhoti2023fine).
-Furthermore, a trivial solution that fine-tunes on the mixture of multilingual and English-only data may be sub-par (e.g., due to language relatedness (dhamecha2021role)).
-Hence, it is challenging to improve an LLM’s language ability while preserving its performance on English.
-Our Approach.
-We use a two-phase continual fine-tuning (CFT) technique for language adaption.
-We study the effects of various English and multilingual instruction tuning datasets when the models are fine-tuned in two phases: where Phase 1 is fine-tuning the model in English to improve its task ability and then fine-tuning it on a proportionally-sized multilingual dataset in Phase 2.
-In Phase 1, we use Alpaca (alpaca) and OpenOrca (OpenOrca), and in Phase 2 we use MultiAlpaca (multialpaca) and mOpenOrca (§3.1).
-We perform this study on two open-source models, namely LLaMA-3-8B and Mistral-7B.
-We also use fine-tuned versions of them, namely LLaMA-3-8B-Instruct and Mistral-7B-Instruct, as off-the-shelf Phase 1 fine-tuned models.
-We quantify a model’s task ability based on its performance on four English datasets: (i) two for instruction following (i.e., IFEval (zhou2023instruction) and Alpaca Eval (li2023alpacaeval)) and (ii) two for reasoning tasks (i.e., MMLU (hendryckstest2021) and HellaSwag (zellers2019hellaswag)).
-Likewise, we quantify a model’s language ability based on its performance on (i) two question answering tasks (i.e., MLQA (lewis2019mlqa) and XQuAD (Artetxe:etal:2019)) and (ii) XLSUM (hasan-etal-2021-xl), a summarization task.
-Our Contributions.
-First, we observe that when phase-wise English and multilingual datasets encode different tasks, we see a decline in the Phase 2 model’s performance on English.
-On the other hand, when Phase 1 and Phase 2 datasets encode similar tasks, the Phase 2 model’s performance on English improves (refer to Figure 1).
-Second, to quantify the similarity of these phase-wise datasets, we introduce two metrics based on language-agnostic embeddings and model representations.
-We show that our quantification correlates with the decline in task ability (§3.3).
-Third, we study the efficacy of two tailored variants of existing CFT strategies to mitigate the decline in task ability after Phase 2 fine-tuning, while also boosting the language ability.
-The first strategy we study is generative replay, i.e., using instructions from a similar English counterpart of the Phase 2 dataset to generate replay data using the Phase 1 model.
-The second strategy uses heuristic-based layer freezing.
-Here, we use the weight difference between the Base and Phase 1 models to pick specific layers for freezing during Phase 2 fine-tuning.
-We study the gains in task and language ability of these strategies compared to specific baselines (§4).
-2 Enhancing Language Ability through Continual Fine-tuning
-A common recipe to training LLMs to learn new languages is to use a training paradigm that focuses on task and language adaption (chen2024improvinglanguageplasticitypretraining).
-Concretely, we define task adaption as the model’s ability to comprehend the input text and then provide a suitable output.
-We refer to language adaption as the model’s ability to perform those tasks in languages other than English.
-In Task Adaption, the LLM is trained to follow instructions, usually using labeled English data.
-Language Adaption focuses on training the LLM to understand text from newer languages.
-Task adaptation leverages cross-lingual transfer, facilitating language adaptation to a certain degree.
-However, this process can result in a decline in the LLM’s task adaptation performance due to the risk of catastrophic forgetting.
-Despite this challenge, task adaptation often yields greater benefits compared to relying solely on cross-lingual transfer for language adaptation.
-Continual Fine-tuning for Language Adaption.
-To improve the language adaption of LLMs, we re-imagine the above recipe as a two-phase CFT process. We have:
-  • Phase 1. We fine-tune a base LLM end-to-end on an English instruction dataset. Phase 1 aims to predominantly teach the LLM instruction following ability, which we refer to as task ability.
-  • Phase 2.
-Here, we use the fine-tuned LLM from Phase 1 and further end-to-end fine-tune it on a Multilingual instruction dataset.
-Unlike chen2024improvinglanguageplasticitypretraining, in our setting, the data in Phase 2 is labeled.
-However, compared to Phase 1, Phase 2’s dataset is geared towards enhancing the LLM’s language ability, and comprises multiple languages with fewer data points per language.
-This paper relates English fine-tuning with task ability enhancement as English fine-tuning predominantly helps in the task ability of LLMs.
-Whereas multilingual fine-tuning predominantly helps with an LLM’s language ability.
-CFT for Language Adaption: Challenges.
-The primary challenge in our two-phase fine-tuning process is that the LLM’s language ability must not come at the cost of its task ability.
+Figure 1: Comparing t-SNEs (van der Maaten and Hinton, 2008) of the hidden activations for Mistral-7B and LLaMA-3-8B during our two-phase Continual Fine-tuning (CFT) process.
+We prompt each model with examples from MTBench Zheng et al.
+(2024), and visualize the similarity between the mean hidden activations, for each model layer.
+For datasets that encode "similar" instructions (Alpaca & MultiAlpaca), English ability does not decline (e.g., 3% gain for IFEval).
+For non-similar datasets (Instruct & MultiAlpaca), English ability declines (e.g., 8% decline for IFEval).
+Here, Phase 2 model representations do not align with Phase 1’s; thus, suggesting greater model weight interference and a decline in English ability.
+The widespread adoption of Large Language Models (LLMs) has led to a growing multilingual user base Shiyas (2023).
+However, ensuring strong performance across languages remains a fundamental challenge, with models consistently performing worse on low-resource languages spoken by millions of speakers worldwide (Ahuja et al., 2023, 2024a).
+A key limitation is that both labeled and unlabeled training data are predominantly available in English and a few high-resource languages, while resources for other languages, especially low-resource ones, are scarce Shaham et al.
+(2024).
+Training large models from scratch is computationally expensive, making fine-tuning pre-trained LLMs the preferred approach for improving multilingual capabilities Lankford et al.
+(2023); Nguyen et al.
+(2023).
+A common fine-tuning strategy is to train LLMs on an instruction-following dataset that contains a mixture of languages.
+However, these datasets are often heavily skewed toward English and other high-resource languages, leading to a performance imbalance: models perform strongly in English but struggle with low-resource languages Dhamecha et al.
+(2021); Li et al.
+(2024a, b).
+Further, prior works show that fine-tuning on a dataset that only contains non-English languages can hurt the model’s performance on English due to catastrophic forgetting Wu et al.
+(2025); Behrouz et al.
+(2025), which is not desirable for most real-world scenarios due to the volume of English queries Ta (2023).
+Ideally, we want the same model to be proficient in both English and other languages to avoid the costs of maintaining multiple models.
+We refer to an LLM’s proficiency in English as its English Ability (EA), and its effectiveness across other languages its Multilingual Ability (MA).
+In this work, we aim to improve an LLM’s MA while maintaining or improving its EA.
+1.1 Our Approach
+To bridge the gap between EA and MA, we introduce a two-phase Continual Fine-tuning (CFT) setup.
+We fine-tune a pre-trained LLM on an English instruction dataset in Phase 1 and then fine-tune it on a similarly-sized Multilingual dataset in Phase 2.
+In Phase 1, we use Alpaca Taori et al.
+(2023) and OpenOrca Lian et al.
+(2023), and in Phase 2 we use MultiAlpaca Wei et al.
+(2023) and mOpenOrca (§4.1).
+Alpaca and OpenOrca provide high-quality English instruction data, while MultiAlpaca and mOpenOrca are their multilingual counterparts, ensuring consistency in instruction style across phases.
+To compare the efficacy of our two-phase CFT setup, we compare it with a straightforward single-phase setup where the LLM is fine-tuned on the mixture of both the instruction tuning datasets.
+We focus on two open-source models, LLaMA-3-8B and Mistral-7B as base models for our experiments.
+We also use fine-tuned versions of them, LLaMA-3-8B-Instruct and Mistral-7B-Instruct, as off-the-shelf Phase 1 English fine-tuned models [^1] LLaMA-3-8B’s pre-training data was 5% multilingual, but LLaMA-3-8B-Instruct is primarily non-multilingual (Dubey et al., 2024)..
+We quantify a model’s English Ability (EA) based on its performance on four English datasets: (i) Two datasets that measure instruction following capabilities (i.e., IFEval Zhou et al.
+(2023) and Alpaca Eval Li et al.
+(2023)) and (ii) two that measure reasoning abilities (i.e., MMLU Hendrycks et al.
+(2021) and HellaSwag Zellers et al.
+(2019)).
+Likewise, we quantify a model’s Multilingual Ability (MA) based on its performance on (i) two question-answering tasks (i.e., MLQA Lewis et al.
+(2019) and XQuAD Artetxe et al.
+(2019)) and (ii) XLSUM Hasan et al.
+(2021), a summarization task.
+1.2 Our Contributions
+CFT Outperforms Mixture.
+We first observe that models trained using our two-phase CFT setup perform better than the single-phase "dataset mixture" setup (Tables 12; §4.2).
+Moreover, our two-phase CFT setup overall results in a better model for all languages, including English, for the same number of training steps.
+The two-phase CFT pipeline also provides more flexibility than training on a mixture of datasets, with the possibility of extending our approach to multi-phase fine-tuning, especially when data from earlier phases might not be available.
+Forgetting vs. Dataset Similarity.
+As mentioned earlier, fine-tuning with multilingual datasets to enhance a model’s multilingual ability can lead to a decline in its English ability due to catastrophic forgetting (Mukhoti et al., 2023; Winata et al., 2023).
+We investigate the factors that may lead to such forgetting by computing the similarity of English and Multilingual Instruction Fine-tuning (IFT) datasets.
+We observe that when English and multilingual datasets have instructions that are not similar, there is a decline in the Phase 2 model’s performance in English.
+On the other hand, when Phase 1 and Phase 2 datasets encode similar instructions, the Phase 2 model’s performance in English improves (refer to Figure 1).
+To quantify the similarity of these phase-wise datasets, we introduce two metrics based on language-agnostic embeddings and model representations.
+We show that our quantification correlates with the decline in English ability (Tables 34; §4.3).
+Mitigating Forgetting.
+We study the efficacy of two tailored variants of existing CFT strategies to mitigate the decline in EA after Phase 2 fine-tuning, while boosting MA.
+The first strategy is distribution replay.
+Here, we look at generative replay, i.e., using instructions from a similar English counterpart of the Phase 2 dataset to generate replay data using the Phase 1 model.
+We also try english replay which acts as language replay by utilizing existing English parallel data from the Phase 2 distribution.
+The second strategy employs layer freezing.
+Our heuristic selects specific layers for freezing during Phase 2 fine-tuning based on the weight differences between the Base and Phase 1 models.
+We also explore Spectrum Hartford et al.
+(2024) as an alternative heuristic.
+We study the gains in EA and MA of these strategies compared to specific baselines (Table 5; §5).
+To the best of our knowledge, we are the first to explore the effectiveness of CFT on LLMs with multilingual instruction datasets.
+2 Related Work
+Continual Learning in LLMs Chen et al. (2026).
+In general, continual learning in LLMs can be broadly categorized into (i) continual pre-training (CPT) and (ii) continual fine-tuning (CFT).
+In CPT, the LLMs are continuously pre-trained to adapt to new domains or tasks by continuously updating them with new data alongside the existing data Shi et al.
+(2024).
+CPT builds on the existing LLM’s knowledge and is more computationally efficient than retraining an LLM using the current and old pre-training data Gupta et al.
+(2023).
+CPT is employed when distributional shifts occur (i) over time Amba Hombaiah et al.
+(2021); Jang et al.
+(2022a, b), (ii) across languages Jin et al.
+(2022); Fujii et al.
+(2024); Blevins et al.
+(2024) or (iii) across domains Ke et al.
+(2023); Gong et al.
+(2022); Xie et al.
+(2023).
+On the other hand, CFT involves training the LLM on successive downstream tasks with varying data distribution or time shifts Shi et al.
+(2024).
+CFT comprises fine-tuning for different tasks Carrión and Casacuberta (2022); Guan et al.
+(2025), instruction-tuning Cahyawijaya et al.
+(2023); Kang et al.
+(2025), model refinement/editing Zhang et al.
+(2023) and alignment Suhr and Artzi (2023).
+Recent literature also focuses on using CFT to assist the LLM to learn new languages Praharaj and Matveeva (2023); Pfeiffer et al.
+(2022); Badola et al.
+(2023); Singh et al.
+(2024c).
+CFT: Enhancing LLMs Multilingual Abilities.
+Cahyawijaya et al.
+Cahyawijaya et al.
+(2023) propose InstructAlign which uses cross-lingual alignment and episodic replay to align an LLM’s pre-trained languages to unseen languages but requires parallel data and previous task data.
+Shaham et al.
+Shaham et al.
+(2024) introduces multilinguality during the first instruction fine-tuning phase which improves an LLM’s instruction following capability across languages.
+He et al.
+He et al.
+(2023) show catastrophic forgetting during CFT and use techniques such as joint fine-tuning and model regularization to mitigate it.
+However, these techniques are computationally expensive or require access to previous task data.
+Multilingual Adaptation.
+This set of works looks at language and task adaption by adjusting the model to understand new languages and enhancing its performance on specific tasks through fine-tuning, respectively Chen et al.
+(2023); Zhao et al.
+(2024); Pfeiffer et al.
+(2020).
+For instance, Chen et al.
+(2023) perform task adaption by fine-tuning the model on downstream task data.
+For language adaption, they fine-tune only the token embedding layer, helping the model learn specific lexical meanings of new languages.
+Language and english ability are either trained in parallel or sequentially.
+However, in this paper, we try to incorporate multilingual ability in models with the constraint that they may have already learned english ability (e.g., Mistral-7B-Instruct).
+To the best of our knowledge, this is a first attempt at studying the effect of task and language self-instruct datasets on an LLM’s multilingual ability through CFT.
+3 Two-phase Continual Fine-tuning Setup
+When instruction fine-tuning LLMs, the most natural method is to fine-tune on a "dataset mixture" containing English and Multilingual data Workshop et al.
+(2023).
+However, fine-tuning on all languages simultaneously may introduce performance bias where the model performs better in English (and other high resource languages) Dhamecha et al.
+(2021); Li et al.
+(2024a, b) [^2] In §4.2, we compare dataset mixture to CFT..
+Continual Fine-tuning (CFT).
+To improve the multilingual performance of pre-trained LLMs, we introduce the following two-phase CFT process.
+Two-Phase CFT Process Phase 1: Fine-tune a base LLM end-to-end on an English instruction dataset.
+Phase 1 aims to teach the LLM English Instruction Following Ability, which we refer to as English Ability (EA).
+Phase 2: Take the fine-tuned LLM from Phase 1 and further fine-tune it end-to-end on a Multilingual instruction dataset.
+Phase 2 focuses on enhancing the LLM’s Multilingual Ability (MA), using a dataset with multiple languages and fewer data points per language.
+Challenges.
+The primary challenge in our two-phase CFT process is that the LLM’s Multilingual Ability must not come at the cost of its English Ability.
 We impose two additional constraints based on real-world scenarios.
 First, in Phase 2, we cannot re-use Phase 1’s dataset.
-Often instruction fine-tuned LLMs are available without their corresponding datasets (e.g., Mistral-7B-Instruct (mistral)).
+Often instruction fine-tuned LLMs are available without their corresponding datasets (e.g., Mistral-7B-Instruct Jiang et al.
+(2023)).
 Second, in Phase 2, we cannot use the weights of the Phase 1 model during training, as saving both old and new set of parameters on the GPU for training would be computationally expensive.
-In a nutshell, we focus on CFT for language adaption for an LLM while preserving the model’s task ability.
-Model Phase 1 (P1) Phase 2 (P2) IFEval ()($\uparrow$)( ↑ ) Alpaca Eval ()($\uparrow$)( ↑ ) MMLU ()($\uparrow$)( ↑ ) HellaSwag ()($\uparrow$)( ↑ ) Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca MultiAlpaca 0.364 0.395 0.12 0.16 0.552 0.573 0.581 0.616 0.404 0.436 Instruct 0.550 0.462 0.35 0.15 0.575 0.533 0.641 0.416 0.529 0.390 LLaMA-3-8B Alpaca 0.277 0.326 0.10 0.11 0.231 0.242 0.556 0.567 0.291 0.311 Instruct 0.735 0.182 0.14 0.10 0.340 0.239 0.533 0.278 0.437 0.2
-Table 1: Task Ability results for two-phase Continual Fine-tuning (CFT).
-When the phase-wise datasets are similar (Definition 1 and Definition 2), task ability post Phase 2 (P2) fine-tuning consistently improves (denoted with green).
-When the phase-wise datasets are not similar, we see a significant decline in task ability post Phase 2 (P2) fine-tuning (denote with red).
-3 Evaluating Task & Language Ability for Multilingual CFT
+Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca 0.364 0.395 0.12 0.16 0.552 0.573 0.581 0.616 0.10 0.11 0.343 0.371 Instruct Multi 0.550 0.462 0.35 0.15 0.575 0.533 0.641 0.416 0.13 0.10 0.449 0.332 LLaMA-3-8B Alpaca Alpaca 0.277 0.326 0.10 0.11 0.231 0.242 0.556 0.567 0.07 0.08 0.247 0.265 Instruct 0.735 0.182 0.14 0.10 0.340 0.239 0.533 0.278 0.11 0.09 0.372 0.178 Dataset Mixture Model Dataset Mixture IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Mistral-7B Alpaca MultiAlpaca 0.394 0.23 0.538 0.602 0.09 0.371 LLaMA-3-8B Alpaca MultiAlpaca 0.363 0.07 0.598 0.602 0.04 0.335
+Table 1: English Ability results for two-phase Continual Fine-tuning (CFT).
+When the phase-wise datasets are similar (Definition 1 and Definition 2), English Ability post Phase 2 (P2) fine-tuning consistently improves (denoted with green).
+When the phase-wise datasets are not similar, we see a significant decline in English Ability post Phase 2 (P2) fine-tuning (denote with red).
+We also provide numbers for dataset mixture – when the models are fine-tuned simultaneously on the Phase 1 and Phase 2 datasets.
+4 Evaluating English & Multilingual Ability for Multilingual CFT
 4.1 Experiment Setup & Evaluation Tasks
 Fine-tuning Models.
-We continually fine-tune open-source Mistral-7B (mistral) and LLaMA-3-8B (llama) LLMs for language adaption.
+We continually fine-tune open-source Mistral-7B Jiang et al. (2023) and LLaMA-3-8B Dubey et al. (2024) LLMs for multilingual adaptation.
 Fine-tuning Datasets.
-For our phase-wise datasets, we use the open-source Alpaca (alpaca), MultiAlpaca (multialpaca), and OpenOrca (OpenOrca) datasets.
+For our phase-wise datasets, we use the open-source Alpaca Taori et al.
+(2023), MultiAlpaca Wei et al.
+(2023), and OpenOrca Lian et al.
+(2023) datasets.
 Alpaca is a self-instruct English-only dataset.
@@
 Fine-tuning Technique.
-We perform full fine-tuning with bf16 precision to study the effects of full fine-tuning with multilingual data in Phase 2 and its effect on task ability.
-We also wish to exploit the benefits gained via complete fine-tuning of these models, which may not be possible with parameter efficient fine-tuning (aggarwal2024maple; panda2024lotteryticketadaptationmitigating).
-However, in §4, we propose a heuristic-based layer freezing strategy to mitigate forgetting of task ability in which we freeze some layers and fine-tune the rest.
-For our experiments, we use Axolotl222github.com/axolotl-ai-cloud/axolotl/, an open-source framework to fine-tune LLMs.
-We conducted our experiments on NVIDIA A100 GPUs with 80 GB RAM.
+We perform full fine-tuning in bf16 precision to study the impact of multilingual post-training in Phase 2 on English ability.
+This setting allows us to fully leverage model capacity, which may not be attainable under parameter-efficient fine-tuning methods Aggarwal et al.
+(2024); Panda et al.
+(2024).
+To mitigate the resulting degradation in English performance, §5 introduces a heuristic-based layer freezing strategy that selectively freezes a subset of layers while fine-tuning the remainder.
+All experiments are conducted using Axolotl333github.com/axolotl-ai-cloud/axolotl/, an open-source framework for large-scale LLM fine-tuning.
 Evaluation Tasks.
-To quantify an LLM’s task ability, we evaluate Phase 1 and Phase 2 models on two instruction-following tasks (i) IFEval (zhou2023instruction) and (ii) Alpaca Eval (li2023alpacaeval), (iii) MMLU (hendryckstest2021) for problem-solving and (iv) HellaSwag (zellers2019hellaswag) for commonsense reasoning ability.
-To quantify an LLM’s language ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering (MLQA (lewis2019mlqa) & XQuAD (Artetxe:etal:2019)) and summarization (XLSUM (hasan-etal-2021-xl)).
-Further details on these tasks are available in the Appendix (§A.3).
-To evaluate our models on TA and LA, we use LM-Evaluation-Harness333github.com/EleutherAI/lm-evaluation-harness, which is a unified framework for zero/few-shot evaluations of LLMs.
-For both task and language ability, we use zero-shot evaluation.
-For additional details on the training setup, code, and evaluation tasks, we refer the reader to the Appendix (§A).
-3.2 Task and Language Ability Results
-Model Phase 1 Phase 2 MLQA ()($\uparrow$)( ↑ ) XLSUM ()($\uparrow$)( ↑ ) XQuAD ()($\uparrow$)( ↑ ) Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B Alpaca MultiAlpaca 0.229 0.288 0.012 0.060 0.290 0.602 0.177 0.317 Instruct 0.246 0.307 0.012 0.033 0.351 0.436 0.203 0.259 LLaMA-3-8B Alpaca 0.438 0.597 0.033 0.034 0.586 0.737 0.352 0.456 Instruct 0.609 0.321 0.048 0.027 0.712 0.417 0.456 0.255
-Table 2: Language Ability results for two-phase Continual Fine-tuning (CFT).
-With green, we denote an improvement in language ability post Phase 2 fine-tuning.
-Likewise, we denote a decline in language ability with red.
-For MLQA and XQUAD we use F1 abstractive score, while for XLSUM we use ROUGE Score.
-We compare the task and language ability of Mistral-7B and LLaMA-3-8B continually fine-tuned models on different phase-wise datasets [^4] When it is clear from the context, we use “Instruct” to denote the dataset used in Phase 1 to instruction fine-tune Mistral-7B-Instruct or LLaMA-3-8B-Instruct..
-Table 1 presents the results for task ability, while Table 2 presents the results for language ability.
+To quantify an LLM’s english ability, we evaluate Phase 1 and Phase 2 models on two instruction-following tasks (i) IFEval Zhou et al.
+(2023) and (ii) Alpaca Eval Li et al.
+(2023), (iii) MMLU Hendrycks et al.
+(2021) for problem-solving, (iv) HellaSwag Zellers et al.
+(2019) for commonsense reasoning ability, and (v) XLSUM_en Hasan et al.
+(2021) for summarization.
+To quantify an LLM’s multilingual ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering (MLQA Lewis et al.
+(2019) & XQuAD Artetxe et al.
+(2019)), instruction-following (GMMLU Singh et al.
+(2024a)), and summarization (XLSUM Hasan et al.
+(2021)).
+Our evaluation suite spans diverse linguistic phenomena and reasoning skills to comprehensively assess language understanding, generation, and problem-solving.
+It also incorporates parallel benchmarks for English and multilingual ability, enabling direct cross-lingual comparison.
+Further details are provided in §A.3.
+To evaluate our models on EA and MA, we use LM-Evaluation-Harness444github.com/EleutherAI/lm-evaluation-harness, which is a unified framework for zero/few-shot evaluations of LLMs.
+For both English and multilingual ability, we use zero-shot evaluation.
+For additional details on the training setup, code, and evaluation tasks, refer to §A.
+4.2 Results
+Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca 0.229 0.288 0.012 0.060 0.290 0.602 0.42 0.40 0.238 0.338 Instruct Multi 0.246 0.307 0.012 0.033 0.351 0.436 0.44 0.43 0.262 0.302 LLaMA-3-8B Alpaca Alpaca 0.438 0.597 0.033 0.034 0.586 0.737 0.53 0.34 0.397 0.427 Instruct 0.609 0.321 0.048 0.027 0.712 0.417 0.25 0.44 0.405 0.301 Dataset Mixture Model Dataset Mixture MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Mistral-7B Alpaca MultiAlpaca 0.406 0.079 0.217 0.41 0.278 LLaMA-3-8B Alpaca MultiAlpaca 0.480 0.040 0.139 0.50 0.289
+Table 2: Multilingual Ability results for two-phase Continual Fine-tuning (CFT).
+With green, we denote an improvement in Multilingual Ability post Phase 2 fine-tuning.
+Likewise, we denote a decline in Multilingual Ability with red.
+For MLQA and XQUAD we use F1 abstractive score, while for XLSUM we use ROUGE Lin (2004) score.
+We also provide numbers for dataset mixture – when the models are fine-tuned simultaneously on the Phase 1 and Phase 2 datasets.
+We compare the English and Multilingual ability of Mistral-7B and LLaMA-3-8B continually fine-tuned models on different phase-wise datasets [^5] When it is clear from the context, we use “Instruct” to denote the dataset used in Phase 1 to instruction fine-tune Mistral-7B-Instruct or LLaMA-3-8B-Instruct..
+Table 1 presents the results for English Ability (EA), while Table 2 presents the results for Multilingual Ability (MA).
 Table 2 reports the average score across languages.
-We provide language-specific scores in the Appendix (§B).
-Results Discussion.
-From Table 1, we see that for phase-wise datasets like Instruct and MultiAlpaca, the performance of the Phase 2 models trained on them declines for English.
+We provide language-specific scores and results when the phases are reversed (e.g., MultiAlpaca-Alpaca) in §B and §C.
+Comparison with Mixture.
+From Tables 1 & 2: for Mixture, the mean of EA and MA scores for Mistral-7B fine-tuned on Alpaca-MultiAlpaca is 0.325, and 0.312 for LLaMA-3-8B.
+The corresponding two-phase mean score is 0.355 for Mistral-7B and 0.346 for LLaMA-3-8B.
+Our two-phase CFT setup is more effective than Mixture, for approximately the same number of training steps.
+Discussion.
+From Table 1, for phase-wise datasets like Instruct and MultiAlpaca, the performance of the Phase 2 models trained on them declines for English.
 This decline occurs when they are continually fine-tuned on multilingual data in Phase 2.
-However, we see a jump in Mistral-7B’s language ability from the results for the multilingual generative tasks (Table 2).
-These models fine-tuned on multilingual datasets show catastrophic forgetting in English.
-However, for phase-wise datasets like Alpaca followed by MultiAlpaca, we see that models trained on them do not show a decline in task ability (Table 1).
-We also see a gain in these models’ language ability (Table 2) [^5] LLaMA-3-8B Instruct MultiAlpaca shows deterioration in LA.
-We explain this behavior in §4.3..
-Additional Ablations.
-In the Appendix (§B), we also present results for OpenOrca-mOpenOrca phase-wise datasets.
-For Mistral-7B, we observe that the average task ability of the Phase 2 model (over Phase 1’s Mistral-7B-OpenOrca) marginally declines: 0.487 from 0.504.
-Whereas, for Mistral-7B-Instruct, the average decline in task ability is significant: 0.376 from 0.529.
-Likewise, for LLaMA-3-8B, the average task ability for LLaMA-3-8B OpenOrca mOpenOrca sees an increase of 0.415 from 0.404.
-In contrast, with Instruct-mOpenOrca as the phase-wise datasets, the task ability significantly drops, from 0.437 to 0.173.
+However, we see a jump in Mistral-7B’s multilingual ability for the multilingual generative tasks (Table 2).
+That is, Phase 2 models fine-tuned on multilingual datasets show forgetting in English.
+However, for phase-wise datasets like Alpaca followed by MultiAlpaca, we see that Phase 2 models do not show a decline in English ability (Table 1).
+We also see a gain in these models’ multilingual ability (Table 2).
+Ablations.
+In Tables B2 & B3 (§B), we present results for OpenOrca-mOpenOrca phase-wise datasets.
+First, the "dataset mixture" again performs worse on average than CFT: 0.186 vs.
+0.372 for Mistral-7B and 0.217 vs.
+0.366 for LLaMA-3-8B.
+Second, for Mistral-7B, the average English ability of the Phase 2 model (over Phase 1’s Mistral-7B-OpenOrca) marginally improves: 0.413 from 0.407.
+Whereas, for Mistral-7B-Instruct, the average decline in English ability is significant: 0.30 from 0.450.
+Likewise, for LLaMA-3-8B, the average English ability for LLaMA-3-8B OpenOrca mOpenOrca sees an increase to 0.356 from 0.337.
+In contrast, for Instruct-mOpenOrca, the English ability significantly drops, from 0.372 to 0.138.
 Observation.
-With Table 1, we see that our two-phase CFT setup for language adaption shows an interesting trend: for certain pairs of phase-wise datasets (e.g., Alpaca & MultiAlpaca), the LLM after Phase 2 sees an improvement in the task ability (computed on English evaluation tasks).
+With Table 1, we see that our two-phase CFT setup for multilingual adaptation shows an interesting trend: for certain pairs of phase-wise datasets (e.g., Alpaca & MultiAlpaca), the LLM after Phase 2 sees an improvement in the English ability (computed on English evaluation tasks).
 We notice that phase-wise datasets like Alpaca and MultiAlpaca have the same seed prompts.
-Alternately, the two datasets encode the same tasks in different languages.
-We hypothesize an LLM fine-tuned on either of these datasets learns the same task ability, and therefore, the second phase of CFT leads to lesser interference in the representation space.
-That is, an LLM continually fine-tuned on Alpaca & MultiAlpaca preserves its task ability across phases.
-We next define two metrics that aim to quantify the task-specific similarity of two datasets.
-3.3 Phase-wise Datasets: Similarity of Representations
+Alternately, the two datasets encode the same instructions in different languages.
+We hypothesize an LLM fine-tuned on either of these datasets learns the same instructions, and therefore, the second phase of CFT leads to lesser interference in the representation space.
+That is, an LLM continually fine-tuned on Alpaca & MultiAlpaca preserves its English ability across phases.
+We next define two metrics that aim to quantify the instruction-specific similarity of two datasets.
+4.3 Similarity of Phase-wise Datasets
+Phase 1 Dataset Phase 2 Dataset DES ($\uparrow$) Alpaca MultiAlpaca 0.924 mOpenOrca 0.792 OpenOrca mOpenOrca 0.953 MultiAlpaca 0.774 Mistral-7B Instruct MultiAlpaca 0.746 : Prepared using model responses on MTBench Zheng et al.
+(2024)
+Table 3: Quantifying Phase-wise Dataset Similarity using DES: higher the score, greater the dataset similarity.
 Dataset Embedding Similarity (DES).
-To quantify whether two datasets encode the same tasks, we define DES that computes a similarity score using the dot product of the average representations (embedding) generated by a language-agnostic model.
+To quantify whether two datasets are similar [^6] The CL-ML literature often defines task similarity via permutation tasks, emphasizing input-output transformations Goldfarb et al.
+(2024).
+Whereas, we consider semantic and structural similarity in natural language instructions., we define DES that computes a similarity score using the dot product of the average representations (embeddings) generated by a language-agnostic model.
 Definition 1 (Dataset Embedding Similarity (DES)).
 Given a language-agnostic text embedding model $\Theta$, and any pair of datasets $D_{1}$ and $D_{2}$, let DES be the function $f_{\textsf{DES}}:D\times D\rightarrow[0,1]$
-$f_{\textsf{DES}}(D_{1},D_{2};\Theta)=\langle\mathbf{E}_{\Theta}(D_{1}),\mathbf{E}_{\Theta}(D_{2})\rangle$ (1)
+$f_{\textsf{DES}}(D_{1},D_{2};\Theta)=\langle\mathbf{E}_{\Theta}(D_{1}),\mathbf{E}_{\Theta}(D_{2})\rangle$
 Here, $\mathbf{E}_{\Theta}(D_{i})\in\mathbb{R}^{d},~\forall i\in\{1,2\}$ is the normalized mean embedding across samples in $D_{i}$.
 Higher the DES score, more similar the embedding, i.e., greater similarity between $D_{1}$ and $D_{2}$.
-For $\Theta$, we use the language-agnostic sentence-tokenizer LaBSE (feng2020languageagnostic).
+For $\Theta$, we use the language-agnostic sentence-tokenizer LaBSE Feng et al.
+(2020).
 We compute DES by encoding 500 random samples from Alpaca, MultiAlpaca, OpenOrca, and mOpenOrca, and measure $f_{\textsf{DES}}$ for each pair.
-Fixing Alpaca as the Phase 1 dataset $D_{1}$, when the Phase 2 dataset $D_{2}$ is MultiAlpaca, the DES score is 0.924 and 0.792 for mOpenOrca.
-When $D_{1}$ is OpenOrca, the DES score for mOpenOrca as $D_{2}$ is 0.953 and 0.774 when $D_{2}$ is MultiAlpaca.
-For dataset pairs with similar tasks, we see a high DES score and relatively low scores for datasets with different tasks.
-That is, DES captures the (pair-wise) variation in task abilities of these datasets.
+Table 3 presents the numbers.
+For dataset pairs with similar datasets, we see a high DES score and relatively low scores for dissimilar datasets.
+DES captures the (pair-wise) variation in instruction similarity of these datasets.
+Dataset $D_{2}$ Model Parameter Difference ($\downarrow$) Alpaca 0.29 Instruct 1.00 OpenOrca 0.55
+Table 4: Quantifying Phase-wise Similarity using MPD: lower the score, greater the dataset similarity. Here, we fix MultiAlpaca as $D_{1}$ and $\theta_{B}$ as Mistral-7B.
 Model Parameter Difference (MPD).
-Another method of quantifying the similarity of tasks for two datasets $D_{1}$ and $D_{2}$ is to compute the difference between the parameters of models $\Theta_{1}$ (fine-tuned on $D_{1}$) and $\Theta_{2}$ (fine-tuned on $D_{2}$).
+Another method of quantifying the similarity of instructions for two datasets $D_{1}$ and $D_{2}$ is to compute the difference between the parameters of models $\Theta_{1}$ (fine-tuned on $D_{1}$) and $\Theta_{2}$ (fine-tuned on $D_{2}$).
 Geometrically, the difference of the parameters captures the representation shift of $\Theta_{2}$ in the space defined by $\Theta_{1}$.
-If $D_{1}$ $D_{2}$ encode the same tasks, the combined shift by $\Theta_{2}$ should be relatively lower, compared to the shift if $D_{1}$ $D_{2}$ encode different tasks.
+If $D_{1}$ & $D_{2}$ encode the same datasets, the combined shift by $\Theta_{2}$ should be relatively lower, compared to the shift if $D_{1}$ & $D_{2}$ encode different intstructions.
 Formally,
@@
 Given any two models $\Theta_{1}$ and $\Theta_{2}$ fine-tuned on self-instruct datasets $D_{1}$ and $D_{2}$ respectively, from the same base model $\Theta_{B}$, let MPD be the function $f_{\textsf{MPD}}:\Theta\times\Theta\rightarrow\mathbb{R}_{\geq 0}$ s.t.
-$f_{\textsf{MPD}}(\Theta_{1},\Theta_{2};\Theta_{B})=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{w}(\Theta_{1,i})-\mathbf{w}(\Theta_{2,i})\|_{2}$ (2)
-[…]
+$f_{\textsf{MPD}}(\Theta_{1},\Theta_{2};\Theta_{B})=\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{w}(\Theta_{1,i})-\mathbf{w}(\Theta_{2,i})\|_{2}$
+Here, $\mathbf{w}(\Theta_{j,i}),~\forall j\in\{1,2\}$ is $\Theta_{j}$’s $i\textsuperscript{th}$ parameter.
 The smaller the MPD score, the closer the fine-tuned models are in the parameter space.
 Fixing Mistral-7B as the base model $\Theta_{B}$, and $D_{1}$ as MultiAlpaca, we vary $D_{2}$ as one of AlpacaOpenOrca, and mOpenOrca, and observe the corresponding MPD scores.
-We normalize the MPD scores with the maximum observed score across all three models for a fair comparison.
-With $D_{2}$ as Alpaca, the MPD score is $0.294$.
-For $D_{2}$ as Instruct, MPD is $1.0$ and 0.55 when $D_{2}$ is OpenOrca.
-These scores show a similar trend to DES: for Alpacaand MultiAlpaca, the scores are lower, highlighting the similarities in the datasets in the parameter space.
+We normalize the MPD scores with the maximum observed score across all three models for a fair comparison (see Table 4).
+MPD shows a similar trend to DES: for Alpaca and MultiAlpaca, the scores are lower, highlighting the similarities in the datasets in the parameter space.
 We see relatively higher scores for the other pair of models, implying a difference in the dataset pairs.
-3.4 Visualizing Decline in Task Ability
+4.4 Visualizing Decline in English Ability
 Setup.
-To explain the effect of similar phase-wise data sets on the LLM task ability, we look at the model representations when parsing English (as the task ability is computed over English).
-We feed MTBench (zheng2024judging) to the models, an English prompt dataset for testing, and visualize the similarity between the mean hidden activations, for each model layer.
+To explain the effect of similar phase-wise data sets on an LLM’s EA, we look at model representations when parsing English.
+We feed MTBench Zheng et al.
+(2024) to the models, a widely-used English benchmark for generalized instruction-following evaluation, and visualize the similarity between the mean hidden activations for each model layer.
 For the analysis, given an LLM $\Theta$ with $l$ layers, let $X_{\Theta}\in\mathbb{R}^{l\times d}$ be the mean hidden activations, across $n$ samples from MTBench.
-Figure 2: We see a greater change in the variation of the representations for non-similar datasets (e.g., Instruct & MultiAlpaca) compared to similar datasets (e.g., Alpaca & MultiAlpaca).
-Interestingly, for LLaMA-3-8B the change is large across layers and a magnitude higher than Mistral-7B.
-For Mistral-7B, we see the later layers showing the most change.
 t-SNE Visualization.
-Figure 1 depicts t-SNEs (vanDerMaaten2008) for $X_{\textsc{Mistral-7B}}$ and $X_{\textsc{LLaMA-3-8B}}$ LLMs, continually fine-tuned on the phase-wise datasets Alpaca & MultiAlpaca and Instruct & MultiAlpaca.
+Figure 1 depicts t-SNEs van der Maaten and Hinton (2008) for $X_{\textsc{Mistral-7B}}$ and $X_{\textsc{LLaMA-3-8B}}$ when these are continually fine-tuned on (i) Alpaca & MultiAlpaca and (ii) Instruct & MultiAlpaca.
 We observe that for similar phase-wise datasets, the model before and after Phase 2 produces similar hidden activations.
@@
 That is, the model representations for non-similar phase-wise datasets are well-separated.
-The model representations during Phase 2 do not align with Phase 1 representations; thus, resulting in greater model weight interference leading to a decline in task ability.
+The separation between model representations results in increased weight interference during Phase 2 – leading to a decline in EA.
 Visualizing Variance in Model Representations.
-Figure 1 provides some intuition for the correlation between phase-wise datasets and decline in task ability.
-To further understand the layer-wise behavior of the hidden activations, similar to chang-etal-2022-geometry, we compute covariance matrices $\Sigma_{\Theta}$ for each $X_{\Theta}$.
+Figure 1 provides an intuition for the correlation between phase-wise datasets and the decline in English ability.
+To further understand the layer-wise behavior of the hidden activations, similar to Chang et al.
+(2022), we compute covariance matrices $\Sigma_{\Theta}$ for each $X_{\Theta}$.
 Intuitively, $\Sigma_{\Theta}$ captures the variance in different directions for representations of hidden activations for $\Theta$.
-We first compute the mean centered activation matrix $\bar{X}_{\Theta}=X_{\Theta}-\mu_{\Theta}$, according to $\mu_{\Theta}\in\mathbb{R}^{d}$.
+Figure 2: Visualizing variance in model representations.
+We plot $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$ across layers to measure representation shift.
+Dissimilar datasets (e.g., Instruct & MultiAlpaca) induce substantially larger variation than similar ones (e.g., Alpaca& MultiAlpaca).
+For LLaMA-3-8B, this variation is consistently high across layers, whereas for Mistral-7B, it is more localized and increases primarily in higher layers.
+We first compute the mean centered activation matrix $\bar{X}_{\Theta}=X_{\Theta}-\mu_{\Theta}$, where $\mu_{\Theta}=1/l\sum_{i=1}^{l}X_{\Theta}^{(i)}$.
 Next, we derive $\Sigma_{\Theta}=\frac{1}{l-1}\cdot\bar{X}_{\Theta}^{T}\bar{X}_{\Theta}\in\mathbb{R}^{d\times d}$.
@@
 This gap is significantly larger for LLaMA-3-8B.
-CFT Setup Task Ability (TA) Language Ability (LA) Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Dataset Strategy ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) ()($\uparrow$)( ↑ ) Mistral-7B 0.55 0.35 0.575 0.641 0.529 0.246 0.012 0.351 0.203 MultiAlpaca 0.462 0.15 0.533 0.416 0.390 0.307 0.033 0.436 0.259 LF_H1 0.456 0.03 0.497 0.598 0.395 0.176 0.016 0.215 0.136 LF_H2 0.364 0.12 0.364 0.504 0.338 0.213 0.014 0.442 0.223 GR_5 0.540 0.17 0.540 0.611 0.465 0.311 0.008 0.428 0.249 GR_10 0.567 0.12 0.567 0.594 0.462 0.213 0.007 0.427 0.215 LoRA 0.383 0.09 0.579 0.625 0.42 0.289 0.043 0.518 0.283 ER_10 0.593 0.08 0.580 0.635 0.599 0.249 0.008 0.398 0.218 LLaMA-3-8B 0.735 0.14 0.340 0.533 0.436 0.609 0.048 0.712 0.456 MultiAlpaca 0.182 0.10 0.239 0.278 0.217 0.321 0.030 0.417 0.256 LF_H1 0.303 0.0 0.231 0.275 0.202 0.368 0.037 0.505 0.303 LF_H2 0.380 0.06 0.485 0.525 0.373 0.400 0.038 0.505 0.314 GR_5 0.269 0.01 0.516 0.316 0.279 0.437 0.019 0.593 0.349 GR_10 0.264 0.12 0.229 0.250 0.228 0.254 0.009 0.314 0.192 LoRA 0.196 0.0 0.280 0.235 0.179 0.007 0.008 0.005 0.007 ER_10 0.420 0.02 0.603 0.561 0.420 0.434 0.025 0.53 0.330
-Table 3: Task and Language ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10) and Layer Freezing (LF_H1 & LF_H2).
-We also use LoRA (hu2022lora) and ER_10 as two baseline strategies.
-Here, we perform Phase 2 fine-tuning with rank 64 and quantisation bfloat16 for LoRA.
-For ER_10, we use the English dataset used in GR_5 with original responses.
-The Phase 1 dataset is Instruct for each row.
-The first two rows for both Mistral-7B and LLaMA-3-8B provide numbers for Instruct and Instruct-MultiAlpaca (from Table 1 & Table 2).
-4 Mitigating Strategies for Multilingual CFT
-To mitigate the decline in task ability, we study two CFT techniques, Generative Replay (GR) and heuristic-based Layer Freezing (LF).
-In Generative Replay, we consider a new English data generation method motivated by the correlation between dataset similarity and task ability (§3.2).
-With heuristic-based Layer Freezing, we employ specific heuristics to find out the subset of layers to freeze in the model during Phase 2 fine-tuning.
-4.1 Generative Replay
-Typically, Generative Replay (GR) is a technique that generates data from past distributions to be used alongside new task data for the continual fine-tuning of a model on a new task (shin2017continuallearningdeepgenerative).
-However, from §3.2, we see that if the phase-wise datasets encode similar tasks, the decline in task ability is mitigated.
-Based on this observation, we use the Phase 1 model to generate responses, in English, from the English counterpart of the multilingual dataset used for training in Phase 2.
-This generated replay dataset acts as a bridge between the distributions of Phase 1 and Phase 2.
+CFT Setup English Ability (EA) Multilingual Ability (MA) Combined Mitigating IFEval Alpaca Eval MMLU HellaSwag XLSUM_en Avg MLQA XLSUM XQUAD GMMLU Avg Avg Strategy $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ $(\uparrow)$ Mistral-7B 0.462 0.15 0.533 0.416 0.10 0.332 0.307 0.033 0.436 0.430 0.302 0.317 LF_H1 0.456 0.03 0.497 0.598 0.10 0.336 0.176 0.016 0.215 0.428 0.209 0.273 LF_H2 0.364 0.12 0.364 0.504 0.12 0.294 0.213 0.014 0.442 0.384 0.263 0.279 Spectrum 0.435 0.24 0.488 0.524 0.13 0.363 0.317 0.083 0.176 0.370 0.237 0.30 GR_5 0.540 0.17 0.540 0.611 0.11 0.394 0.311 0.008 0.428 0.445 0.298 0.346 GR_10 0.567 0.12 0.567 0.594 0.12 0.394 0.213 0.007 0.427 0.450 0.274 0.334 ER_10 0.593 0.08 0.580 0.635 0.13 0.404 0.249 0.008 0.398 0.448 0.276 0.340 LoRA 0.383 0.09 0.579 0.625 0.03 0.341 0.289 0.043 0.518 0.435 0.321 0.331 LLaMA-3-8B 0.182 0.10 0.239 0.278 0.09 0.178 0.321 0.030 0.417 0.440 0.302 0.240 LF_H1 0.303 0.0 0.231 0.275 0.01 0.164 0.368 0.037 0.505 0.237 0.287 0.225 LF_H2 0.380 0.06 0.485 0.525 0.08 0.306 0.400 0.038 0.505 0.338 0.320 0.313 Spectrum 0.409 0.09 0.612 0.524 0.01 0.329 0.429 0.056 0.086 0.472 0.261 0.295 GR_5 0.269 0.01 0.516 0.316 0.07 0.236 0.437 0.019 0.593 0.342 0.348 0.292 GR_10 0.264 0.12 0.229 0.250 0.0 0.173 0.254 0.009 0.314 0.238 0.204 0.189 ER_10 0.420 0.02 0.603 0.561 0.12 0.345 0.434 0.025 0.53 0.448 0.359 0.352 LoRA 0.196 0.0 0.280 0.235 0.0 0.142 0.007 0.008 0.005 0.278 0.075 0.109
+Table 5: English Ability (EA) and Multilingual Ability (MA) results for our mitigating strategies.
+These comprise Generative Replay (GR_5 & GR_10), English Replay (ER_10) and Layer Freezing (LF_H1, LF_H2 & Spectrum).
+We use LoRA Hu et al.
+(2022) as a baseline strategy.
+For ER_10, we use the English dataset used in GR with original responses.
+The Phase 1 dataset is Instruct for each row, while Phase 2 is MultiAlpaca.
+The first row for both Mistral-7B and LLaMA-3-8B provides numbers for Instruct-MultiAlpaca (from Table 1 &  2).
+5 Mitigating Strategies for CFT
+To mitigate EA decline, we explore two tailored CFT techniques [^7] In §F, we discuss how MAD-X Pfeiffer et al.
+(2020) and other PEFT-based methods Badola et al.
+(2023) are not suitable for our setting.: Distribution Replay and Layer Freezing.
+In Distribution Replay, we study Generative Replay (GR), a new English data generation method inspired by dataset similarity and English ability (§4.2), and English Replay (ER), which replays parallel English data of Phase 2’s distribution.
+In Layer Freezing (LF), we identify layers to freeze during Phase 2 fine-tuning using specific heuristics.
+5.1 Distribution Replay
+Typically, Generative Replay (GR) is a technique that generates data from past distributions to be used alongside new task data for the continual fine-tuning of a model on a new task Shin et al.
+(2017).
+However, from §4.2, we do not see a decline in English ability if the phase-wise datasets encode similar instructions.
+Based on this, we use the Phase 1 model to generate responses, in English, from the English counterpart of the multilingual dataset used for fine-tuning in Phase 2.
+The intuition is that the generated dataset may bridge the distributions of Phase 1 and Phase 2.
 During Phase 2 fine-tuning, we include varying quantities of this generated data: specifically, 5% (GR_5) and 10% (GR_10), of the Phase 2 dataset.
-As a baseline, we also fine-tune the models with a similar sized subset of the English counterpart with original responses [^6] This dataset may not be available for all multilingual datasets eg.
-Aya (singh2024ayadatasetopenaccesscollection).
-In that case, instructions can always be translated to English but it is not always practical to translate responses.
-Hence, this baseline is the best-case scenario for our GR strategy..
-We refer to this baseline as English Replay (ER_10).
-4.2 Heuristic-based Layer Freezing
-Model regularization is an effective technique to mitigate the drop in the previous task’s performance in continual learning (e.g., EWC (kirkpatrick2017overcoming)).
+We also fine-tune the models with a similar sized subset of the English counterpart with original responses [^8] This dataset may not be available for all multilingual datasets, such as Aya Singh et al.
+(2024b).
+While instructions can be translated into English, translating responses is often impractical.
+Thus, ER is the best-case scenario for GR..
+We refer to this mitigating strategy as English Replay (ER_10).
+5.2 Layer Freezing
+Model regularization is an effective technique to mitigate the drop in the previous task’s performance in continual learning (e.g., EWC Kirkpatrick et al.
+(2017)).
 However, this is computationally inefficient as it requires using both the old and new sets of parameters.
-Instead, we use Layer Freezing (LF), a relatively efficient technique for use as a ‘regularizer’ to preserve task ability during Phase 2.
-We consider the following two variations to select the set of layers to freeze:
+Instead, we use Layer Freezing (LF), a relatively efficient technique for use as a ‘regularizer’ to preserve English ability during Phase 2.
+We consider the following variations to select the set of layers to freeze.
+These variants allow us to study how different criteria for selecting frozen layers affect the trade-off between preserving English ability and enabling multilingual adaptation.
   1. 1.
@@
     LF_H2: freezing the top-10 layers that have changed the most during Phase 1 fine-tuning (e.g., Mistral-7B Base to Mistral-7B-Instruct).
-We select these layers separately for Key, Query, and Value, for each attention head.
-We present our results in Table 3 for both GR and LF. Along with English Replay (ER), we define another baseline in which we use LoRA (hu2022lora) for continually fine-tuning in Phase 2.
+We select layers separately for Key, Query, and Value, for each attention head.
+  3. 3.
+    Spectrum Hartford et al. (2024): freeze the "most informative" layers of the Phase 1 model based on their signal-to-noise ratio (refer to §D.1).
+We present our results in Table 5 for both GR and LF.
+We define a baseline in which we use LoRA Hu et al.
+(2022) [^9] Parameter efficient techniques like LoRA Hu et al.
+(2022) are also widely used to efficiently fine-tune LLMs on multilingual data.
+However, such techniques also show forgetting on English Aggarwal et al.
+(2024) after Phase 2.for continually fine-tuning in Phase 2.
+We perform LoRA fine-tuning with rank 64 and quantisation bfloat16.
 5.3 Results Discussion
-From Table 3, we see that GR and LF successfully mitigate the decline in task ability and also show gains in language ability.
-For instance, Mistral-7B with GR_5 achieves better performance in MLQA and XLSUM when fine-tuned with MultiAlpaca.
-We also close the gap with Mistral-7B-Instruct  on IFEval, Alpaca Eval, MMLU, and HellaSwag with our mitigation strategies.
-(a) Mistral-7B
-(b) LLaMA-3-8B
-Figure 3: Demonstrating extent of cross-lingual transfer in Mistral-7B and LLaMA-3-8B on a parallel dataset prepared by subsampling Flores (costa2022no).
-We find that the English activation cluster for LLaMA-3-8B is separated from the multilingual cluster, compared to Mistral-7B.
-LLaMA-3-8B Doesn’t Show Consistent Improvement with our Mitigations.
-From Table 3, while both GR and LF improve on the baseline LLaMA-3-8B-Instruct MultiAlpaca, the gains in task and language ability are not comparable to LLaMA-3-8B-Instruct.
-To understand this further, for GR, we investigate the cross-linguality difference between LLaMA-3-8B and Mistral-7B.
-Like Figure 1, we plot t-SNEs of the mean model activations for the Mistral-7B and LLaMA-3-8B base models on two parallel datasets, English and Multilingual.
-We create the parallel datasets by subsampling data from Flores (costa2022no).
-In Figure 3, we see that the English activation cluster for LLaMA-3-8B is separated out from multilingual cluster, compared to Mistral-7B.
-This suggests that GR may not be as effective when the model has less cross lingual ability.
-While for LF, we acknowledge that our method to identify the layers to freeze may not be the best and better methods to identify which layers to freeze can be a direction for future work.
-Last, but not the least, we acknowledge that LLaMA-3-8B-Instruct seems to be a strong model even on multilingual benchmarks.
-Hence, it is also important to evaluate Phase 1 models on these benchmarks first and then decide if the Phase 2 fine-tuning step should be undertaken or not.
-With regards to LLaMA-3-8B-Instruct MultiAlpaca LA results in Table 2, we believe that this is due to lack of cross-linguality in LLaMA-3-8B-Instruct and less data in MultiAlpaca which fails to cause sufficient representation drift to improve the model’s performance.
-Forgetting with LoRA.
-For Mistral-7B-Instruct and LoRA fine-tuning, we see an increase in language ability but a decline in task ability.
-But the decline is not as much as full fine-tuning.
-For LLaMA-3-8B-Instruct and LoRA, there is a greater decline in both task and language ability.
-The decline is similar (or slightly lower) than the full fine-tuning scenario.
-These results show that LoRA also suffers from forgetting when used for continual fine-tuning.
-Additional Results.
-In the Appendix (§C), we repeat the same experiment from §3.4 to quantify the representation change in the fine-tuned models using our mitigating strategies.
-We see a trend similar to Figure 2.
-That is, a decrease in the variation in the model activations, compared to the baseline model trained on Instruct and MultiAlpaca.
-The decrease is more pronounced for Mistral-7B compared to LLaMA-3-8B.
-In Appendix §C, we also present TA and LA results for the Instruct-mOpenOrca phase-wise datasets.
+From Table 5, we see that GR, ER, and LF mitigate the decline in EA and also show gains in MA.
+Distribution Replay.
+ER_10 demonstrates the best performance in both English and combined ability, with EA scores of 0.404 for Mistral-7B and 0.345 for LLaMA-3-8B, and the best combined average.
+GR_5 also excels in multilingual tasks, performing similar to ER_10: 0.298 vs.
+0.276 for Mistral-7B and 0.348 vs.
+0.359 for LLaMA-3-8B.
+GR_5 also performs reasonably well on English tasks, achieving scores of 0.394 and 0.236 for Mistral-7B and LLaMA-3-8B, respectively, making it a competitive strategy.
+Layer Freezing.
+Compared to ER and GR, LF_H1, LF_H2, and Spectrum show mixed results.
+LF_H2 performs better than LF_H1.
+Spectrum’s EA scores are better than LF_H1 and LF_H2, but suffers from lower multilingual numbers.
+Additional Discussion & Results.
+In §D.3, we also present EA and MA results for Mistral-7B Instruct-mOpenOrca for our mitigating strategies.
+Here, LF, particularly Spectrum, performs better than the other strategies.
+Furthermore, in §D.4, we analyze the computational cost of these strategies over the baseline CFT setup.
+5.4 Controlling Representation Drift under Mitigation
+Figure 3: Layer-wise covariance shift under mitigation strategies in Phase 2 CFT.
+We plot $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$ across layers.
+The vanilla Phase 2 model (Figure 2) exhibits a sharp increase in drift in higher layers.
+Replay (GR) substantially reduces the magnitude of this drift across all layers, while freezing (LF_H2) confines it to a smaller subset of layers.
+Building on Figure 2, we examine how mitigation strategies affect representation drift relative to the vanilla Phase 2 baseline (Mistral-7B Instruct $\rightarrow$ MultiAlpaca).
+As in §4.4, we measure drift as $\|\Sigma_{\text{Phase 2}}-\Sigma_{\text{Phase 1}}\|_{2}$, where larger values indicate greater deviation from Phase 1 representations.
+Consistent with earlier observations, the mitigation strategies explicitly designed to curb representational change (LF & GR) exhibit lower drift than the baseline model across layers.
+Replay-based methods (GR) reduce the overall magnitude of drift while preserving its layer-wise structure, whereas freezing-based methods (LF_H2) constrain the drift to a subset of layers.
+Notably, generative replay remains closest to the Phase 1 representation geometry among all strategies.
+This closer alignment correlates with improved task and language performance compared to the vanilla Phase 2 model (Table 1 and Table 2).
+These results reinforce our central hypothesis: in two-phase CFT with dissimilar datasets, representation shift is the primary driver of English degradation, and explicitly controlling this shift preserves alignment with Phase 1 representations, leading to improved English retention without sacrificing multilingual gains.
 6 Conclusion & Future Work
-In this paper, to the best of our knowledge, we present a first study on the influence of the similarity of phase-wise datasets on the task and language adaptability of LLMs through CFT.
-Through extensive experiments on the Mistral-7B and LLaMA-3-8B models, we show that when datasets are similar, task ability is preserved; otherwise, it declines.
-To mitigate this, we propose layer freezing and generative replay as mitigating strategies based on specific heuristics.
-Our results indicate that these strategies help improve task performance while not compromising on the LLM’s language adaptability.
+We study the role of dataset similarity in a novel two-phase continual fine-tuning (CFT) for multilingual adaptation.
+Across Mistral-7B and LLaMA-3-8B, we show that alignment between Phase 1 and Phase 2 datasets is critical: similar datasets preserve English ability while improving multilingual performance, whereas misaligned datasets lead to degradation.
+We identify representation shift as the underlying mechanism driving this behavior.
+In particular, misaligned fine-tuning induces large covariance shifts in hidden representations, concentrated in higher layers.
+We show that mitigation strategies such as generative replay and layer freezing improve English retention by controlling this drift, either by globally reducing its magnitude or by constraining it to specific layers, while preserving multilingual gains.
 Future Work.
-Our results show that there is no one-size-fits-all strategy to mitigate decline in task ability, among the strategies discussed.
-Future work can explore developing other parameter-efficient regularization methods that address the current computational challenges with methods like EWC or forgetting due to LoRA.
-One can also explore analytical notions for task similarity in datasets.
+Our results suggest several directions for future research.
+First, designing adaptive mitigation strategies that dynamically control representation drift during training could provide stronger and more efficient trade-offs than fixed heuristics Zhang et al.
+(2026).
+Second, developing principled measures of instruction-level dataset similarity may enable better prediction and prevention of degradation.
+Finally, exploring parameter-efficient approaches that explicitly regularize representation geometry could reduce the computational overhead of current methods while maintaining their effectiveness.
+7 Limitations
+Our study has several limitations.
+First, we rely on DES and MPD as proxies for dataset similarity; while effective in our setting, these metrics may not capture all nuances of instruction-level similarity.
+Second, our experiments are limited to Mistral-7B and LLaMA-3-8B, and the observed trends may not fully generalize to models with different architectures, scales, or training paradigms.
+Similarly, the effectiveness of our mitigation strategies (i.e., replay and layer freezing) may vary across datasets and model families.
+Third, our experimental setup is constrained by computational resources.
+We were unable to explore larger models, longer training schedules, or broader hyperparameter sweeps, which may further influence the observed trade-offs.
+Fourth, some mitigation strategies impose additional requirements.
+In particular, the best-performing method, ER_10, assumes access to parallel data, which may not always be available in practice.
+Finally, our evaluation focuses on specific benchmarks for task and language ability.
+While these provide controlled comparisons, they may not fully capture real-world performance across diverse applications.
 Appendix A Training Details
 A.1 Hyperparameters for Fine-tuning and Training Setup
-Hyperparameter Value Learning Rate $1\times 10^{-6}$ Epochs 4 Global Batch size 16 Scheduler Cosine Warmup Linear Warmup Steps 10 Optimizer AdamW (loshchilov2019decoupled) Weight Decay 0
-Table A1: Hyperparameters for continual fine-tuning
+Hyperparameter Value Learning Rate $1\times 10^{-6}$ Epochs 4 Global Batch size 16 Scheduler Cosine Warmup Linear Warmup Steps 10 Optimizer AdamW Loshchilov and Hutter (2019) Weight Decay 0
+Table A1: Hyperparameters for our Two-phase Continual Fine-tuning (CFT)
 A.2 Fine-tuning Datasets
-OpenOrca is an English-only self instruct dataset, created to best mimic the Orca dataset (mukherjee2023orca), which is not publicly available.
-To create the multilingual version of OpenOrca, namely mOpenOrca, we follow sphinx to generate selective translations for a subset of OpenOrca.
+OpenOrca is an English-only self instruct dataset, created to best mimic the Orca dataset Mukherjee et al.
+(2023), which is not publicly available.
+To create the multilingual version of OpenOrca, namely mOpenOrca, we follow Ahuja et al.
+(2024b) to generate selective translations for a subset of OpenOrca.
 The subset contains 50k samples from the OpenOrca dataset and we selectively translate them to 11 languages which are also in MultiAlpaca.
@@
 In this paper, we consider two sets of benchmarks to evaluate task and language ability. We explain them briefly next.
-Task Ability (TA).
+English Ability (EA).
 To quantify an LLM’s task ability, we evaluate Phase 1 and Phase 2 models on the following tasks:
   1. 1.
-    IFEval (zhou2023instruction): Instruction-Following Evaluation (IFEval) asses the ability of an LLM to follow natural language instructions.
+    IFEval Zhou et al.
+(2023): Instruction-Following Evaluation (IFEval) asses the ability of an LLM to follow natural language instructions.
 It comprises 500 verifiable instructions (e.g., “mention the keyword AI 3 times”).
@@
   2. 2.
-    Alpaca Eval (li2023alpacaeval): This is an LLM-based automatic evaluator for instruction following models, to measure task ability.
-Like aggarwal2024maple, we evaluate our CFT models against text-davinci-003 responses on 800 instructions and use GPT4 (gpt-4-32k) as the evaluator.
+    Alpaca Eval Li et al.
+(2023): This is an LLM-based automatic evaluator for instruction following models, to measure task ability.
+Like Aggarwal et al.
+(2024), we evaluate our CFT models against text-davinci-003 responses on 800 instructions and use GPT4 (gpt-4-32k) as the evaluator.
   3. 3.
-    MMLU (hendryckstest2021): Massive Multitask Language Understanding (MMLU) is a benchmark to assess an LLM’s knowledge and problem-solving abilities.
+    MMLU Hendrycks et al.
+(2021): Massive Multitask Language Understanding (MMLU) is a benchmark to assess an LLM’s knowledge and problem-solving abilities.
 It includes 57 subjects across domains like STEM, or law, with 16k MCQs in total.
   4. 4.
-    HellaSwag (zellers2019hellaswag): This is a popular benchmark to evaluate the commonsense reasoning ability of an LLM. HellaSwag’s test split contains 10k samples in total.
-Language Ability (LA).
+    HellaSwag Zellers et al. (2019): This is a popular benchmark to evaluate the commonsense reasoning ability of an LLM. HellaSwag’s test split contains 10k samples in total.
+Multilingual Ability (MA).
 To quantify an LLM’s language ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering and summarization.
-  • Question Answering: MLQA (lewis2019mlqa) contains 5k extractive question-answering instances in 7 languages.
-The XQuAD dataset (Artetxe:etal:2019) consists of a subset of 240 paragraphs and 1190 question-answer pairs across 11 languages.
-  • Summarisation: XLSUM (hasan-etal-2021-xl) spans 45 languages, and we evaluate our models in Arabic, Chinese-Simplified, English, French, Hindi, Japanese, and Spanish.
-Model Phase 1 (P1) Phase 2 (P2) IFEval ()($\uparrow$)( ↑ ) Alpaca Eval ()($\uparrow$)( ↑ ) MMLU ()($\uparrow$)( ↑ ) HellaSwag ()($\uparrow$)( ↑ ) Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.494 0.482 0.31 0.32 0.601 0.582 0.612 0.562 0.504 0.487 Instruct 0.550 0.426 0.35 0.06 0.575 0.507 0.641 0.509 0.529 0.376 LLaMA-3-8B OpenOrca 0.377 0.425 0.09 0.07 0.579 0.599 0.571 0.564 0.404 0.415 Instruct 0.735 0.205 0.14 0.0 0.340 0.236 0.533 0.250 0.437 0.173
-Table A1: Task Ability results for two-phase Continual Fine-tuning (CFT).
+  • Question Answering: MLQA Lewis et al.
+(2019) contains 5k extractive question-answering instances in 7 languages.
+The XQuAD dataset Artetxe et al.
+(2019) consists of a subset of 240 paragraphs and 1190 question-answer pairs across $1$ languages.
+  • Summarisation: XLSUM Hasan et al. (2021) spans 45 languages, and we evaluate our models in Arabic, Chinese-Simplified, English, French, Hindi, Japanese, and Spanish.
+Appendix B Evaluating Multilingual Ability for Continual Fine-tuning
+Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.494 0.482 0.31 0.32 0.601 0.582 0.612 0.562 0.02 0.12 0.407 0.413 Instruct mOpenOrca 0.550 0.426 0.35 0.06 0.575 0.507 0.641 0.509 0.13 0.0 0.450 0.30 LLaMA-3-8B OpenOrca mOpenOrca 0.377 0.425 0.09 0.07 0.579 0.599 0.571 0.564 0.07 0.12 0.337 0.356 Instruct mOpenOrca 0.735 0.205 0.14 0.0 0.340 0.236 0.533 0.250 0.11 0.0 0.372 0.138 Dataset Mixture Model Dataset Mixture IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Mistral-7B OpenOrca mOpenOrca 0.228 0.035 0.284 0.444 0.02 0.202 LLaMA-3-8B OpenOrca mOpenOrca 0.248 0.072 0.484 0.473 0.0 0.255
+Table B2: English Ability results for two-phase Continual Fine-tuning (CFT).
 With green, we highlight an increase in a model’s task ability post P2 fine-tuning.
 Likewise, red highlights a decline in a model’s task ability.
-Model Phase 1 Phase 2 MLQA ()($\uparrow$)( ↑ ) XLSUM ()($\uparrow$)( ↑ ) XQuAD ()($\uparrow$)( ↑ ) Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B OpenOrca mOpenOrca 0.435 0.36 0.007 0.008 0.556 0.643 0.332 0.337 Instruct 0.246 0.155 0.012 0.040 0.351 0.323 0.203 0.173 LLaMA-3-8B OpenOrca 0.401 0.453 0.017 0.006 0.499 0.531 0.306 0.330 Instruct 0.609 0.604 0.048 0.048 0.712 0.713 0.456 0.455
-Table A2: Language Ability results for two-phase Continual Fine-tuning (CFT).
-With green, we highlight an increase in a model’s language ability post Phase 2 fine-tuning.
-Likewise, red highlights a decline in a model’s language ability.
-Appendix B Evaluating Language Ability for Multilingual Continual Fine-tuning
-Task Ability.
-Table A1 present the task ability numbers of our ablations on the OpenOrca-mOpenOrcaand Instruct-mOpenOrcadatasets using Mistral-7B and LLaMA-3-8B models.
-When the datasets are pairwise not similar, i.e., Instruct-mOpenOrca, Mistral-7B shows a significant decline in the average task ability, from 0.529 in Phase 1 to 0.376 in Phase 2.
-Likewise, LLaMA-3-8B also experiences a decrease, dropping from 0.437 to 0.173 on average.
-In contrast, when the pairwise datasets are similar, i.e., OpenOrca and mOpenOrca, Mistral-7B sees a marginal drop between the phases $(0.504\rightarrow 0.487)$, on average.
-LLaMA-3-8B’s performance sees an improvement in the average task ability, from 0.404 to 0.415.
-Language Ability.
-Table A2 tabulates the results for language ability.
-We see an improvement in the average language ability for the OpenOrca-mOpenOrca dataset pair, for both Mistral-7B and LLaMA-3-8B.
-For Instruct-mOpenOrca, with LLaMA-3-8B, the average language ability is virtually the same across tasks.
-However, for Mistral-7B, we see a slight drop in the average language ability, driven primarily due to a drop in performance for MLQA.
+Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B OpenOrca mOpenOrca 0.435 0.360 0.007 0.008 0.556 0.643 0.433 0.312 0.358 0.331 Instruct mOpenOrca 0.246 0.155 0.012 0.040 0.351 0.323 0.440 0.279 0.262 0.20 LLaMA-3-8B OpenOrca mOpenOrca 0.401 0.453 0.017 0.006 0.499 0.531 0.242 0.513 0.290 0.376 Instruct mOpenOrca 0.609 0.604 0.048 0.048 0.712 0.713 0.250 0.233 0.405 0.40 Dataset Mixture Model Dataset Mixture MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ GMMLU $(\uparrow)$ Average Mistral-7B OpenOrca mOpenOrca 0.201 0.128 0.071 0.277 0.169 LLaMA-3-8B OpenOrca mOpenOrca 0.224 0.034 0.091 0.364 0.178
+Table B3: Multilingual Ability results for two-phase Continual Fine-tuning (CFT).
+With green, we highlight an increase in a model’s Multilingual ability post Phase 2 fine-tuning.
+Likewise, red highlights a decline in a model’s Multilingual ability.
+English Ability.
+Table B2 present the english ability numbers of our ablations on the OpenOrca-mOpenOrcaand Instruct-mOpenOrcadatasets using Mistral-7B and LLaMA-3-8B models.
+When the datasets are pairwise not similar, i.e., Instruct-mOpenOrca, Mistral-7B shows a significant decline in the average English ability, from 0.450 in Phase 1 to 0.30 in Phase 2.
+Likewise, LLaMA-3-8B also experiences a decrease, dropping from 0.372 to 0.138 on average.
+In contrast, when the pairwise datasets are similar, i.e., OpenOrca and mOpenOrca, Mistral-7B sees a marginal increase between the phases $(0.407\rightarrow 0.413)$, on average.
+LLaMA-3-8B’s performance sees an improvement in the average English ability, from 0.337 to 0.356.
+Multilingual Ability.
+Table B3 tabulates the results for multilingual ability.
+We see an improvement in the average multilingual ability for the OpenOrca-mOpenOrca dataset pair for LLaMA-3-8B.
+For Mistral-7B, there is a marginal drop $(0.358\to 0.331)$.
+For Instruct-mOpenOrca, with LLaMA-3-8B, the average multilingual ability is virtually the same across tasks (0.405 vs.
+0.40).
+However, for Mistral-7B, we see a slight drop in the average language ability, driven primarily due to a decline in performance for MLQA.
+Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B MultiAlpaca Alpaca 0.245 0.290 0.120 0.114 0.528 0.430 0.476 0.510 0.342 0.336 LLaMA-3-8B MultiAlpaca Alpaca 0.245 0.340 0.038 0.065 0.570 0.540 0.577 0.590 0.357 0.384 Mistral-7B mOpenOrca OpenOrca 0.190 0.310 0.091 0.055 0.410 0.490 0.520 0.510 0.303 0.341 LLaMA-3-8B mOpenOrca OpenOrca 0.314 0.340 0.0 0.0 0.530 0.540 0.522 0.590 0.342 0.368
+Table B4: English Ability results for two-phase Continual Fine-tuning (CFT)
+Model Phase 1 Phase 2 MLQA $(\uparrow)$ XLSUM $(\uparrow)$ XQuAD $(\uparrow)$ Average Dataset Dataset Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Phase 1 Phase 2 Mistral-7B MultiAlpaca Alpaca 0.122 0.230 0.021 0.030 0.122 0.090 0.088 0.116 LLaMA-3-8B MultiAlpaca Alpaca 0.363 0.340 0.048 0.040 0.058 0.030 0.157 0.134 Mistral-7B mOpenOrca OpenOrca 0.165 0.160 0.077 0.070 0.140 0.180 0.127 0.137 LLaMA-3-8B mOpenOrca OpenOrca 0.057 0.0 0.038 0.0 0.047 0.0 0.047 0.0
+Table B5: Multilingual Ability results for two-phase Continual Fine-tuning (CFT)
 Furthermore, Table B6, Table B7, and Table B8 present the language-specific results for MLQA, XLSUM, and XQuAD, respectively.
@@
 Table B8: XQuAD: Language Ability results for two-phase Continual Fine-tuning (CFT).
-Appendix C Mitigating Strategies
-Here, we first visualize the impact of our mitigating strategies on the variance in model representations. Next, we provide ablate our findings for the Instruct-mOpenOrca phase-wise datasets.
-Figure C4: Visualizing Variance in Model Representations for Mistral-7B Mitigating Strategies: We see a decrease in the variance of model representations for models trained using our mitigation strategies compared to vanilla Phase 2 models (see Figure 2).
-Visualizing Variance in Model Representations.
-In Figure C4, we repeat the same experiment as in § 4.5 to quantify the representation change in the fine-tuned models using our mitigating strategies.
-The trend seen is expected from §4.5: we see a decrease in the variation in the model activations, compared to the baseline model trained on Instruct and MultiAlpaca.
-For the mitigating strategies that are curated to curb representational change, i.e., LF_H2, GR_5, and GR_10, we see that the corresponding curves have lesser change than the baseline Phase 2 model, Mistral-7B Instruct MultiAlpaca.
-That is, there is less representational change for LF_H2, GR_5, and GR_10 compared to Mistral-7B Instruct MultiAlpaca.
-Our generative replay techniques are the closest in the representational change to Mistral-7B Instruct.
-This ‘closeness’ also improves its task and language ability performance compared to the vanilla Phase 2 model, Mistral-7B Instruct MultiAlpaca(refer to Table 1 and Table 2).
-Additional Ablations.
-We also present the impact of our mitigating strategies for the Instruct-mOpenOrca phase-wise datasets on Mistral-7B. Table C6 presents these results.
+Appendix C Reverse Order CFT Result Analysis
+In Tables B4 and B5, we reverse the order of Phase 1 and Phase 2 datasets, first fine-tuning on the multilingual dataset followed by its English counterpart.
+For Mistral-7B (MultiAlpaca-Alpaca), the average performance is 0.226, and for LLaMA-3-8B (MultiAlpaca-Alpaca), it is 0.259.
+Compared to the mixture setting and the Alpaca-MultiAlpaca configuration (§4), we observe that English ability benefits from multilingual fine-tuning in Phase 1, resulting in performance comparable to the data mixture setting.
+However, fine-tuning on English data in Phase 2 leads to a drastic drop in multilingual ability, yielding worse results than both the mixture setting and the two-phase setup considered in the main paper.
+Appendix D Mitigating Strategies
+Here, we provide additional details on Spectrum Hartford et al.
+(2024).
+We then visualize the impact of our mitigating strategies on the variance in model representations.
+Lastly, we ablate our findings for the Instruct-mOpenOrca phase-wise datasets.
+D.1 Spectrum
+Spectrum Hartford et al.
+(2024) is a layer-freezing technique that optimizes the fine-tuning of LLMs by selecting layers based on their signal-to-noise ratio (SNR).
+We use Spectrum as a heuristic for layer-freezing; that is, the layers identified as "important" by Spectrum are frozen during Phase 2 fine-tuning.
+A layer is important based on its signal-to-noise (SNR) ratio.
+In the following, we elaborate on how Spectrum computes SNR.
+Marchenko-Pastur distribution.
+The Marchenko-Pastur distribution Marchenko and Pastur (1967) is given by:
+$\rho(\lambda)=\frac{1}{2\pi\sigma^{2}}\frac{\sqrt{(\lambda_{+}-\lambda)(\lambda-\lambda_{-})}}{\lambda},$
+where
+$\lambda_{\pm}=\sigma^{2}(1\pm\sqrt{Q})^{2},$
+and $Q=\frac{N}{M}$, with $N$ and $M$ being the dimensions of a random matrix $W$, and $\sigma^{2}$ representing the variance of the entries in $W$.
+SNR.
+Let $W\in\mathbb{R}^{N\times M}$ be the weight matrix of a given layer.
+The empirical spectral density of $W$ is analyzed by comparing its eigenvalue distribution of $1/N\cdot W^{T}W$ against the theoretical Marchenko-Pastur distribution.
+Deviations from this distribution indicate the presence of significant signal components.
+We get,
+$\lambda_{\pm}=\sigma^{2}\left(1\pm\sqrt{\frac{M}{N}}\right)^{2},$
+where $\lambda_{\pm}$ are the largest and smallest eigenvalues and $\sigma$ the standard deviation. This implies the bounds of singular values of $W$ as:
+$\epsilon_{\pm}=\frac{1}{\sqrt{N}}\sigma\left(1\pm\sqrt{\frac{M}{N}}\right)$ (1)
+By evaluating how the singular values of $W$ distribute relative to $\epsilon_{\pm}$, Spectrum assesses the SNR of each layer, as defined next.
+Ratio
+Hartford et al. (2024). Specifically, the SNR value of a weight matrix is,
+$\text{SNR}=\frac{\sum_{k|\sigma_{k}>\epsilon}\sigma_{k}}{\sum_{k|\sigma_{n}<\epsilon}\sigma_{n}}$
+Here, $\epsilon$ separates signal from noisy singular values.
+Layers with singular values significantly exceeding $\epsilon_{+}$ have a high SNR, indicating a substantial presence of informative signal components.
+Measuring the Ratio
+Hartford et al.
+(2024).
+Having defined all ingredients above, Spectrum now computes each layer’s SNRs.
+To do this, it first computes SVD Zhang and Xu (2009) of the the layer’s weight matrix, calculates the SNR and normalizes it by the highest singular value.
+Eq.
+1 gives the noise threshold.
+Now, Spectrum selects layers with higher SNRs, where the number of layers selected is a hyperparameter.
+Similar to Hartford et al.
+(2024), for our experiments, we select the top-50% of layers in each module.
+D.2 LLaMA-3-8B Doesn’t Show Consistent Improvement with our Mitigation Strategies
+From Table 5, while both GR and LF improve on the baseline LLaMA-3-8B-Instruct MultiAlpaca, the gains in task and multilingual ability are not comparable to LLaMA-3-8B-Instruct.
+(a) Mistral-7B
+(b) LLaMA-3-8B
+Figure D4: Demonstrating extent of cross-lingual transfer in Mistral-7B and LLaMA-3-8B on a parallel dataset prepared by subsampling Flores Costa-jussà et al.
+(2022).
+We find that the English activation cluster for LLaMA-3-8B is separated from the multilingual cluster, compared to Mistral-7B.
+To understand this further, for GR, we investigate the cross-linguality difference between LLaMA-3-8B and Mistral-7B.
+Like Figure 1, we plot t-SNEs of the mean model activations for the Mistral-7B and LLaMA-3-8B base models on two parallel datasets, English and Multilingual.
+We create the parallel datasets by subsampling data from Flores Costa-jussà et al.
+(2022).
+In Figure D4, we see that the English activation cluster for LLaMA-3-8B is separated out from multilingual cluster, compared to Mistral-7B.
+This suggests that GR may not be as effective when the model has less cross lingual ability.
+While for LF, we acknowledge that our method to identify the layers to freeze may not be the best and better methods to identify which layers to freeze can be a direction for future work.
+Last, but not the least, we acknowledge that LLaMA-3-8B-Instruct seems to be a strong model even on multilingual benchmarks.
+Hence, it is also important to evaluate Phase 1 models on these benchmarks first and then decide if the Phase 2 fine-tuning step should be undertaken or not.
+With regards to LLaMA-3-8B-Instruct MultiAlpaca MA results in Table 2, we believe that this is due to lack of cross-linguality in LLaMA-3-8B-Instruct and less data in MultiAlpaca which fails to cause sufficient representation drift to improve the model’s performance.
+D.3 Additional Ablations
+We also present the impact of our mitigating strategies for the Instruct-mOpenOrca phase-wise datasets on Mistral-7B. Table D9 presents these results.
 We see that LF_H2 achieves moderate success, especially in maintaining the language ability for MLQA (0.258) and XQUAD (0.527).
@@
 Perhaps LoRA shows the best overall performance among the strategies for maintaining task ability (e.g., IFEval = 0.587, MMLU = 0.567, HellaSwag = 0.591) with reasonable retention of language ability (e.g., XQUAD = 0.354).
-Note.
 These results show that no single strategy is perfect, and future work may need to combine these strategies or develop new approaches to address the balance between task and language ability retention across phases.
-CFT Setup Task Ability Language Ability Model Phase 1 Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Dataset Dataset Strategy Mistral-7B Instruct 0.55 0.35 0.575 0.641 0.529 0.246 0.012 0.351 0.203 mOpenOrca 0.426 0.06 0.507 0.509 0.376 0.155 0.040 0.323 0.173 LF_H2 0.401 0.048 0.518 0.487 0.364 0.258 0.060 0.527 0.282 GR_5 0.281 0.027 0.478 0.495 0.320 0.167 0.042 0.305 0.171 GR_10 0.305 0.013 0.483 0.494 0.324 0.150 0.038 0.238 0.142 LoRA 0.587 0.13 0.567 0.591 0.469 0.167 0.027 0.354 0.183 ER_10 0.367 0.025 0.479 0.493 0.341 0.157 0.042 0.305 0.168
-Table C6: Task and Language ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10) and Layer Freezing (LF_H1 & LF_H2).
-We also use LoRA (hu2022lora) and ER_10 as two baseline strategies.
-Here, we perform Phase 2 with rank 64 and bf16 for LoRA.
-For ER_10, we use the English dataset used in GR_5 with original responses.
-Appendix D Resources Used
-We used 4 NVIDIA A100 GPU (80 GB) with a 96 core AMD CPU to run our inferences. One Finetuning Run with MultiAlpaca took 4 hours while for mOpenOrca it took 12 hours.
-the list of model and the URL with checkpoints available and licenses are listed below:
-  • LLaMA-3-8B : meta-llama/Meta-Llama-3-8B License: llama3
-  • Mistral-7B : https://huggingface.co/mistralai/Mistral-7B-v0.1 License: Apache-2.0
+CFT Setup Task Ability Language Ability Overall Model Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Avg Dataset Strategy Mistral-7B mOpenOrca 0.426 0.060 0.507 0.509 0.376 0.155 0.040 0.323 0.173 0.275 LF_H2 0.401 0.048 0.518 0.487 0.364 0.258 0.060 0.527 0.282 0.323 Spectrum 0.442 0.158 0.508 0.616 0.431 0.387 0.086 0.201 0.225 0.328 GR_5 0.281 0.027 0.478 0.495 0.320 0.167 0.042 0.305 0.171 0.246 GR_10 0.305 0.013 0.483 0.494 0.324 0.150 0.038 0.238 0.142 0.233 ER_10 0.367 0.025 0.479 0.493 0.341 0.157 0.042 0.305 0.168 0.255 LoRA 0.587 0.130 0.567 0.591 0.469 0.167 0.027 0.354 0.183 0.326
+Table D9: English and Multilingual Ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10), English Replay (ER_10) and Layer Freezing (LF_H1, LF_H2 & Spectrum).
+We use LoRA Hu et al.
+(2022) as a baseline strategy.
+For ER_10, we use the English dataset used in GR with original responses.
+The Phase 1 dataset is Instruct for each row.
+The first row provides Mistral-7B numbers for Instruct-mOpenOrca (from Table B2).
+D.4 Compute Analysis
+The computational overhead of replay arises from an increase in the effective training set size .
+In ER_10, each epoch includes the original Phase 2 data along with an additional 10% replay buffer from Phase 1, resulting in a proportional increase in the number of training tokens and hence compute.
+Similarly, GR_5 augments each epoch with 5% generated replay data, leading to a smaller overhead.
+In contrast, layer freezing (LF) significantly reduces computation by updating only a subset of model parameters.
+By freezing 50% of the layers, LF yields a corresponding reduction in training cost while still achieving a reasonable trade-off between English retention and multilingual adaptation.
+Appendix E Resources Used
+We conduct all experiments on 4 NVIDIA A100 GPUs (80GB each) with a 96-core AMD CPU.
+A single fine-tuning run on MultiAlpaca takes approximately 4 hours, while training on mOpenOrca requires around 12 hours.
+The models used in our experiments, along with their corresponding checkpoints and licenses, are listed below:
+Appendix F Existing Mitigating Strategies
+We now discuss why MAD-X Pfeiffer et al. (2020) and other PEFT-based methods Badola et al. (2023) are not well-suited to our setting.
+  • Parameter-Efficient Fine-tuning for Robust Continual Multilingual Learning Badola et al.
+(2023): This work considers a fixed downstream task (text classification) that is incrementally extended to new languages, mitigating forgetting via task- and language-specific adapters.
+Their approach assumes continued access to prior task data and the ability to maintain multiple adapter checkpoints.
+In contrast, our setting involves end-to-end instruction fine-tuning over a heterogeneous mixture of tasks.
+We do not assume access to Phase 1 data during Phase 2, nor do we maintain multiple parameter sets due to the memory overhead of storing additional model weights (§3).
+  • MAD-X Pfeiffer et al.
+(2020): MAD-X introduces bottleneck adapters within encoder-only or encoder–decoder Transformer architectures to enable cross-lingual transfer.
+However, recent work Zhao et al.
+(2024) shows that such adapters do not integrate cleanly with decoder-only LLMs (e.g., Mistral-7B, LLaMA-3-8B), which form the basis of our experiments.
+As our models lack an encoder stack, the routing and merging mechanisms central to MAD-X are not directly applicable.
+  • Preserving Cross-Linguality via Continual Learning Liu et al.
+(2021): This method extends Gradient Episodic Memory by storing exemplars (or their gradients) from previous phases and replaying them during training.
+Our setup (§3) explicitly disallows (i) reusing Phase 1 data and (ii) maintaining Phase 1 representations or weights during Phase 2, as these incur significant memory overhead and may violate practical data-sharing constraints.